### v1

# Parallel computations for Metropolis Markov chains with Picard maps

Abstract
We develop parallel algorithms for simulating zeroth-order (aka gradient-free) Metropolis Markov chains based on the Picard map. For Random Walk Metropolis Markov chains targeting log-concave distributions $\pi$ on $\mathbb{R}^{d}$, our algorithm generates samples close to $\pi$ in $\mathcal{O}(\sqrt{d})$ parallel iterations with $\mathcal{O}(\sqrt{d})$ processors, therefore speeding up the convergence of the corresponding sequential implementation by a factor $\sqrt{d}$. Furthermore, a modification of our algorithm generates samples from an approximate measure $\pi_{\epsilon}$ in $\mathcal{O}(1)$ parallel iterations and $\mathcal{O}(d)$ processors. We empirically assess the performance of the proposed algorithms in high-dimensional regression problems and an epidemic model where the gradient is unavailable. Our algorithms are straightforward to implement and may constitute a useful tool for practitioners seeking to sample from a prescribed distribution $\pi$ using only point-wise evaluations of $\log\pi$ and parallel computing.
Key words— Parallel computing, Metropolis-Hastings, sampling, log-concavity, Gradient-free, Zeroth-order
1 Introduction
1.1 Picard map for parallel-in-time Markov chain simulation
Consider a discrete-time Markov chain $(X_{i})_{i\in\mathbb{N}}$ on $\mathcal{X}\subseteq\mathbb{R}^{d}$ defined as
$X_{i+1}$ $=X_{i}+f(X_{i},W_{i})\,,$ $i=0,1,\dots$ (1)
for some initial position $X_{0}\in\mathcal{X}$, measurable function $f\colon\mathcal{X}\times\mathcal{W}\to\mathbb{R}^{d}$ and random innovations $W_{0},W_{1},\dots\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\nu$ on $\mathcal{W}$. The canonical way to simulate the first $K$ steps of the chain, $X=(X_{0},\dots,X_{K})$, is to go through the recursion in (1) for $i=0,1,\dots,K-1$. This procedure requires $K$ sequential calls to the function $f$. Alternatively, one can define $X$ through the set of equations
$X_{i+1}$ $=X_{0}+\sum_{\ell=0}^{i}f(X_{\ell},W_{\ell})\,,$ $i=0,1,\dots\,.$
Following this representation, one can define the Picard map (named after the celebrated Picard–Lindelöf theorem), $\Phi\colon\mathcal{X}^{K+1}\times\mathcal{W}^{K}\mapsto\mathcal{X}^{K+1}$, which takes as input a trajectory $X\in\mathcal{X}^{K+1}$ and the innovations $W=(W_{0},\dots,W_{K-1})\in\mathcal{W}^{K}$, and returns a new trajectory $X^{\prime}=(X^{\prime}_{0},\dots,X^{\prime}_{K})=\Phi(X,W)$ defined as
$X_{i}^{\prime}$ $=\Phi_{i}(X,W)=\begin{cases}X_{0}&i=0\\ X_{0}+\sum_{\ell=0}^{i-1}f(X_{\ell},W_{\ell})&0<i\leq K\,.\end{cases}$ (2)
Given $W\in\mathcal{W}^{K}$, the map $\Phi$ is deterministic and its fixed point, i.e. the trajectory $X\in\mathcal{X}^{K+1}$ satisfying $X=\Phi(X,W)$, is unique and coincides with the solution to (1).
This motivates computing $X$ as the limit of the recursion $X^{(j)}=\Phi(X^{(j-1)},W)$ for $j=1,2,\dots$, with initialization $X^{(0)}\in\mathcal{X}^{K+1}$. See Figure 1 for an illustration. A canonical choice for initialization is the constant trajectory $X^{(0)}=(X_{0},\dots,X_{0})\in\mathcal{X}^{K+1}$. The computational advantage of the Picard recursion over the classical one defined in (1) is that the $K$ calls to the function $f$ in (2) can be executed in parallel. This can be useful when computing $f$ is expensive.
The Picard recursion is guaranteed to converge to its fixed point in at most $K$ iterations because, by construction, $X^{(j)}$ coincides with the solution of (1) in the first $j$ coordinates. However, since the sequential recursion in (1) also requires $K$ iterations to compute $X$, relevant applications of the Picard recursion are those where $X^{(j)}$ gets close (or exactly equal) to its fixed point for $j\ll K$ iterations, trading the use of parallel computing for a shorter run-time.
Figure 1: Traces on the $(x_{1},x_{2})$-plane of the Picard recursion $X^{(j)}$ for $K=1000$ and $j=1,2,10,13$ (top left - bottom right). The time-index of each sequence is shown with a yellow-red gradient color. The underlying Markov chain is a $d=100$ dimensional Random Walk Metropolis with step-size $\xi=2/\sqrt{d}$ targeting a standard Gaussian distribution. The gray line is the fixed point (i.e. the output of the sequential algorithm). The dashed line corresponds to the part of the trajectory that has converged to its fixed point.
1.2 Zeroth-order Markov chain Monte Carlo
Markov chain Monte Carlo (MCMC) methods are the computational workhorse of many scientific areas, including Bayesian Statistics. These methods consist of simulating a Markov chain $(X_{0},X_{1},\dots)$ whose limiting measure coincides with a probability distribution of interest $\pi$, such as a Bayesian posterior. Then, ergodic averages $\frac{1}{N}\sum_{i=1}^{N}g(X_{i})$ can be used to estimate expectations $\int g(x)\pi(\mathrm{d}x)$ for any integrable function $g$, see e.g. Roberts and Rosenthal (2004). We are interested in the simulation of zeroth-order (aka gradient-free, derivative-free) Markov chains, which only require point-wise evaluations of $\log\pi$ (up to an additive unknown constant). These methods are commonly used when gradient information is unavailable, for either practical, computational, or mathematical reasons (cf. Conn et al., 2009, Chapter 1). This may occur for example in Bayesian applications with black-box likelihood evaluations (e.g. in form of proprietary code, code written by domain experts, or models requiring complex numerical solvers) or when the gradient is not well-defined such as in models with censored data (e.g. the SIR model considered in Section 6), in Pseudo-Marginal MCMC (Andrieu and Roberts, 2009) and in Approximate Bayes Computation (ABC) MCMC setting (Marjoram et al., 2003). See e.g. Conn et al. (2009) and other references on zeroth-order optimization for more discussion about applications of gradient-free methods.
Our objective is to leverage the use of $K>1$ parallel processors for reducing the parallel round complexity of this class of algorithms, defined as the number of point-wise evaluations of $\log\pi$ per parallel processor in order to obtain samples from $\pi$. Under standard log-concavity assumptions on $\pi$, the best known complexity result w.r.t. the number of dimensions $d$ for gradient-free sequential algorithms (i.e. $K=1$) is $\mathcal{O}(d)$ evaluations per sample, and it is achieved by classical Random Walk Metropolis [^1] Throughout the paper, the $\mathcal{O}$ notation ignores constants and logarithmic terms (Dwivedi et al., 2019; Andrieu et al., 2024). Note that this complexity matches the one of gradient-free convex optimization (Nesterov and Spokoiny, 2017).
1.3 Main results
Corollary 2 shows that, given a log-concave target density $\pi$ on $\mathbb{R}^{d}$ and $K=\mathcal{O}(\sqrt{d})$ processors, our proposed Online Picard algorithm applied to a Random Walk Metropolis Markov chain (initialized from a so-called feasible start) generates samples from a measure close to $\pi$ in total variation distance after $\mathcal{O}(\sqrt{d})$ parallel iterations. Thus, the Online Picard algorithm accelerates zeroth-order sequential algorithms by a factor of $\sqrt{d}$ with $K=\mathcal{O}(\sqrt{d})$ parallel processors, achieving optimal speedup (i.e. linear in $K$ up to logarithmic factors) for $K$ up to $\mathcal{O}(\sqrt{d})$. On the contrary, for larger values of $K$, we observe, both theoretically and empirically, a sharp decrease in relative gain. Motivated by this bottleneck and leveraging insights from the analysis of the Online Picard algorithm, we also propose an Approximate Picard algorithm which generates samples from an approximate measure $\pi_{\epsilon}$ after $\mathcal{O}(1)$ parallel iterations with $K=\mathcal{O}(d)$ processors. The error of the invariant measure $\pi_{\epsilon}$ relative to $\pi$ is assessed through numerical simulations, while a rigorous analysis lies beyond the scope of this work. Section 4 provides convergence results also for Picard algorithms applied to Metropolis within Gibbs Markov chains, which we empirically observe to perform better. The empirical performance observed in Section 6 closely matches the theory developed and shows substantial speedups (up to over 100 times) relative to the corresponding sequential algorithms, with highest efficiency obtained by Metropolis within Gibbs schemes. All proofs are provided in the appendix.
These results, combined with the simplicity of our algorithms, offer practitioners promising directions to parallelize computations for Bayesian problems with black-box, expensive models and no access to gradient information.
1.4 Related literature on parallel sampling and Picard maps
Previous attempts to parallelize computations for (zeroth-order) MCMC include running multiple independent chains (Rosenthal, 2000), which is straightforward to implement but does not reduce the burn-in or convergence period, or using pre-fetching (Brockwell, 2006) and multiple-try (Frenkel and Smit, 1996; Liu et al., 2000) algorithms, which however achieve complexity gains of only $\mathcal{O}(\log(K))$ in typical zeroth-order settings (Pozza and Zanella, 2024). The past year has seen a growing interest in parallel sampling based on Picard recursions applied to first-order (i.e. gradient-based) Markov chains for both Monte Carlo methods (Yu and Dalalyana, 2024; Anari et al., 2024) and generative models (Shih et al., 2023; Zhou and Sugiyama, 2024; Chen et al., 2024; Gupta et al., 2024). However, existing analyses focus on the (so-called unadjusted) setting where the function $f$ in (1) is smooth with respect to their first argument, which is not the case for Metropolis-Hastings algorithms – see (8) below. Our work avoids the classical smoothness conditions on $f$ and instead focuses on Metropolis-type Markov chains, where $f$ is piecewise constant with respect to its first argument. This setting leads to a substantially different analysis as well as convergence results for the induced Picard map, as highlighted in Section 2 and Section 3.4 below.
2 Picard algorithms with piecewise constant map $\Phi$
We consider Markov chains that satisfy the following assumption.
Assumption 1.
For every $W\in\mathcal{W}^{K}$, the Picard map $X\mapsto\Phi(X,W)$ is constant in a neighborhood of its fixed point.
Assumption 1 is satisfied by standard zeroth order Metropolis-Hastings kernels described in Section 3.2, which are our main motivation. Other relevant examples satisfying this assumption include Markov chains defined on discrete spaces. Assumption 1 has many implications. For example, it implies that the fixed point of $\Phi$ can be reached exactly after $j<K$ parallel iterations. When the fixed point is reached, the output of the Picard recursion matches that of the sequential counterpart, with no additional bias induced by the Picard recursion. This is in contrast to classical applications of Picard recursions, where stopping the Picard recursion before $K$ parallel iterations introduces bias and requires tuning of an additional error parameter.
2.1 Online Picard algorithm
Assume we seek to simulate $N$ steps of the Markov chain defined in (1), $(X_{0},\dots,X_{N})$, using $K\leq N$ independent processors. In the classical Picard algorithm, one updates the first $K$ steps, $(X_{1},\dots,X_{K})$, with the Picard recursion until a stopping criterion is met, then proceeds with the next $K$ ones, $(X_{K+1},\dots,X_{2K})$, and so on until reaching $X_{N}$. Under Assumption 1, instead, one can monitor the coordinates that have already reached their fixed point while the algorithm runs, and avoid wasting computational power in updating them further, directly allocating available processors to subsequent coordinates. This leads to the following Online Picard algorithm: given a sequence of noise $W=(W_{0},W_{1},\dots)\in\mathcal{W}^{\infty}$ and initialization $X^{(0)}\in\mathcal{X}^{\infty}$, define the $j$-th parallel recursion of the Online Picard algorithm mapping $X^{(j)}\in\mathcal{X}^{\infty}$ to $X^{(j+1)}\in\mathcal{X}^{\infty}$ as
$X_{i}^{(j+1)}$ =$=\begin{cases}X_{i}^{(j)}&i\leq L^{(j)},\\ \Phi_{i-L^{(j)}}\left(X_{L^{(j)}:U^{(j)}}^{(j)},W_{L^{(j)}:U^{(j)}-1}\right)&L^{(j)}<i\leq U^{(j)}\,,\\ X_{U^{(j)}}^{(j+1)}&i>U^{(j)},\end{cases}$ (3)
where $L^{(0)}=0$, $U^{(j)}=L^{(j)}+K$ and
$L^{(j+1)}$ $=\sup\{i\leq U^{(j)}\colon f(X^{(j)}_{\ell},W_{\ell})=f(X^{(j+1)}_{\ell},W_{\ell})\text{ for }0\leq\ell<i\}$
$=\sup\{i\leq U^{(j)}\colon f(X_{\ell},W_{\ell})=f(X^{(j)}_{\ell},W_{\ell})\text{ for }0\leq\ell<i\}$
$=\sup\{i\leq U^{(j)}\colon X_{\ell}=X^{(j)}_{\ell}\text{ for }0\leq\ell\leq i\},\qquad\hbox{for }j\geq 0\,,$ (4)
where the equivalence of the formulations above follows from (2).
The Online Picard algorithm applies the Picard map to the indices $i\in\{L^{(j)},L^{(j)}+1,\dots,U^{(j)}\}$, where $L^{(j)}$ is an increasing sequence which monitors the position of the last index $i$ such that $X^{(j)}_{0:i}$ has reached its fixed point. The algorithm is then stopped as soon as $L^{(j)}\geq N$. While (3) is formally defined as a map on $\mathcal{X}^{\infty}$, the algorithm in practice only operates on $K$ coordinates of the Markov chain, as shown by Proposition 1 below and by the pseudocode provided in Appendix C. As for the Picard recursion in (2), the online recursion in (3) requires $U^{(j)}-L^{(j)}=K$ calls of the function $f$ that can be computed in parallel.
The Online Picard algorithm is always more efficient than the classical Picard algorithm which applies sequentially the Picard recursion of (2) until convergence of each subsequent block $\{X_{i}\colon\ell K<i\leq(\ell+1)K\}$ for $\ell=0,1,\dots$, since it avoids the vacuous updates of $X^{(j)}_{i}$ for $i\leq L^{(j)}$. See Figure 2 for an illustration.
Figure 2: Illustration of the Picard recursion applied sequentially to every block of length $K$ (left grid) vs the Online Picard algorithm (right grid). The color of the $(j,i)$ entry of each grid represents the state of the $i$th step of the Markov chain at the $j$th Picard recursion: red for $f(X^{(j)}_{i},W_{i})=f(X^{(j-1)}_{i},W_{i})$ (correct guess), blue for $f(X^{(j)}_{i},W_{i})\neq f(X^{(j-1)}_{i},W_{i})$. Black for the increments for which no processor has been allocated for computing the function $f$. Here, $K=4$, $N=3K$. Yellow boundary line in correspondence of $L^{(j)}=\sup\{i\leq U^{(j)}\colon X_{\ell}^{(j)}=X_{\ell}^{(j-1)};\ell\leq i\}$.
2.2 Markov structure of Online Picard algorithms
Let $(X^{(j)},W^{(j)},L^{(j)})_{j=0,1,\dots}$ be the sequence of random variables produced by the Online Picard recursion of (3)-(4). Define $\bar{X}^{(j)}=X^{(j)}_{L^{(j)}:U^{(j)}}$ and $\bar{W}^{(j)}=W_{L^{(j)}:U^{(j)}-1}$.
Proposition 1.
The sequence $(\bar{X}^{(j)},\bar{W}^{(j)})_{j=0,1,\dots}$ is a time-homogeneous Markov chain on $(\mathcal{X}^{K+1}\times\mathcal{W}^{K})$.
Proof.
The sequence $(\bar{X}^{(j)},\bar{W}^{(j)})_{j=0,1,\dots}$ can be described by the recursion
$\bar{X}^{(j+1)}_{i}$ =$=\begin{cases}\Phi_{G^{(j)}+i}(\bar{X}^{(j)},\bar{W}^{(j)})&i\leq K-G^{(j)}\,,\\ \Phi_{K}(\bar{X}^{(j)}\,,\bar{W}^{(j)})&i>K-G^{(j)}\,,\\ \end{cases}$ $i=0,1,\dots,K\,,$ (5)
$\bar{W}^{(j+1)}_{i}$ =$=\begin{cases}\bar{W}^{(j)}_{G^{(j)}+i}&i\leq K-G^{(j)}-1\,,\\ \tilde{W}_{i,j}&i>K-G^{(j)}-1\,,\end{cases}$ $i=0,1,\dots,K-1\,,$ (6)
where $\tilde{W}_{i,j},\,i,j\geq 0,$ are i.i.d. random variables with distribution $\nu$ on $\mathcal{W}$,
$G^{(j)}=G(\bar{X}^{(j)},\bar{W}^{(j)})=\sup\{i\leq K\colon\Phi_{\ell}(\bar{X}^{(j)},\bar{W}^{(j)})=\bar{X}^{(j)}_{\ell}\text{ for }0\leq\ell\leq i\}\,,$ (7)
and (5)-(6) define the one-step Markov transition kernel of $(\bar{X}^{(j)},\bar{W}^{(j)})_{j=0,1,\dots}$. ∎
Proposition 1 sheds lights to the underlying (Markov) structure of Online Picard recursion. It can also be used for a memory-wise and computationally efficient implementation of the algorithm that iterates over a vector of length $K$. In the next section, we study the distribution of the random variables $(L^{(j)})_{j\in\mathbb{N}}$ for popular zeroth-order Markov chains under log-concavity, in order to derive parallel round complexity results for the Online Picard algorithm.
3 Complexity of Picard algorithms for zeroth-order Monte Carlo
3.1 Assumption on the target
We study convergence of the Picard map under the following assumption on $\pi$.
Assumption 2 (Smooth and Hessian Lipschitz target).
The target distribution $\pi$ has a Lebesgue density in $\mathcal{X}=\mathbb{R}^{d}$ proportional to $\exp(-V)$ where $V$ is a twice-differentiable, $L$-smooth and $(\gamma L^{3/2}d^{1/2})$-strong Hessian-Lipschitz for some $\gamma\geq 0$, meaning that the Hessian of $V$ satisfies
$-L\mathrm{I}_{d}$ $\preccurlyeq\nabla^{2}V(x)\preccurlyeq L\mathrm{I}_{d}\,,$ $x\in\mathcal{X}$
$-M\|y-x\|$ $\preccurlyeq\nabla^{2}V(y)-\nabla^{2}V(x)\preccurlyeq M\|y-x\|\,,$ $x,\,y\in\mathcal{X}$
for some $M$ with $\mathrm{Tr}(M)\leq\gamma L^{3/2}d^{1/2}$, where $A\preccurlyeq B$ if $A-B$ is positive semi-definite for two symmetric matrices $A,B$.
The strong Hessian-Lipschitz condition implies the more classical Hessian-Lipschitz condition, $\|\nabla^{2}V(y)-\nabla^{2}V(x)\|\leq m^{\star}\|y-x\|$, where $m^{\star}=\lambda_{max}(M)$, while it is weaker than assuming a bounded Frobenius norm of the third derivative, as previously used for analysis of MCMC algorithms (cf. Chen and Gatmiry, 2023). Note also that the analysis of the convergence of $\Phi$ does not require strong convexity of $V$.
3.2 Metropolis-Hastings kernels
In this work, we focus on two popular zeroth-order Metropolis-Hasting Markov chains: Random Walk Metropolis (RWM) (Metropolis et al., 1953) and Metropolis within Gibbs (MwG) (Geman and Geman, 1984). Both satisfy Assumption 1 and can be simulated with the Online Picard algorithm. We start with the RWM case and defer the MwG one to Section 4.
Assumption 3 (RWM).
$X_{0},X_{1},\dots$ is a RWM Markov chain on $\mathcal{X}=\mathbb{R}^{d}$ with step-size $\xi=h/(\sqrt{Ld})$ for some $h>0$, i.e. it follows (1) with $W=(Z,U)$, $U\sim\mathrm{Unif}([0,1])$ uniformly distributed, $Z\sim\mathcal{N}(0,h^{2}I_{d}/(Ld))$ independently of $U$,
$f(x,W)$ $=ZB(x,U,Z)\quad\hbox{with }B(x,u,z)=\mathbf{1}\left(\pi(x+z)/\pi(x)\geq u\right)$ (8)
and $\mathbf{1}(\cdot)$ being the indicator function.
By (4) and (8) combined we obtain
$L^{(j+1)}$ $=\sup\{i\leq U^{(j)}\colon B(X_{\ell},U_{\ell},Z_{\ell})=B(X^{(j)}_{\ell},U_{\ell},Z_{\ell})\text{ for }0\leq\ell<i\}\,,$ $j\geq 0\,,$
which means that, in the RWM case, the Picard convergence is entirely dependent on the probability of ‘correctly guessing’ the $\{0,1\}$-valued random variables $B(X_{i},U_{i},Z_{i})$.
3.3 Main results
Let $(X_{i})_{i=0,1,\dots}$ and $(X^{(j)}_{i})_{i,j=0,1,2,\dots}$ be as in (1) and (3), respectively.
Theorem 1 (Probability of an incorrect guess).
Under Assumptions 2-3, for all $x_{0}\in\mathcal{X}$, $w_{0}\in\mathcal{W}$, we have
$\mathbb{P}(f(X^{(j)}_{i},W_{i})\neq f(X_{i},W_{i})\mid X_{0}=x_{0},W_{0}=w_{0})$ $\leq c_{0}\frac{i}{d}+\delta(d)+2^{-j}$ (9)
for all $0\leq i\leq\min(d,K)$ and $j\in\{0,1,\dots\}$, with $c_{0}=15h^{4}(\sqrt{\frac{2}{\pi}}+\frac{h\gamma}{2})^{2}$, $\delta(d)=\frac{5}{3}\exp(-3d/2)$. The probability in (9) is with respect to the randomness of $(W_{1},W_{2},\dots)$.
Theorem 1 essentially states that, after a logarithmic number of Picard recursions, the probability that the Picard map correctly predicts the $i$th step is $\mathcal{O}(i/d)$.
Remark 1.
By Theorem 1, the probability of correctly guessing individual increments can be controlled up to $i=\mathcal{O}(d)$. Assuming (9) to be tight, this also suggests that increasing the number of parallel processors beyond $K=\mathcal{O}(d)$ provides little gains in terms of speeding-up convergence of the Picard recursion, as numerically observed in Section 6.
Remark 2.
The upper bound of Theorem 1 becomes vacuous for $i>d$. This is consistent with the fact that the mixing time of RWM under Assumptions 2 is $\mathcal{O}(d)$. Indeed, if $X_{0}$ and $X_{i}$ are approximately independent, then the prediction of $B(X_{i},U_{i},Z_{i})$ with $B(X_{0},U_{i},Z_{i})$ is uninformative.
By the union bound, Theorem 1 implies that, with high probability, the Online Picard algorithm makes $\mathcal{O}(\sqrt{d})$ consecutive correct guesses after $\mathcal{O}(\log(d))$ calls of the Picard map when using $K=\mathcal{O}(\sqrt{d})$ processors, as shown below.
Corollary 1.
Under Assumptions 2-3, for all $x_{0}\in\mathcal{X},\,w_{0}\in\mathcal{W}$, $\epsilon>0$, $K\leq\sqrt{\epsilon^{2}d/(2c_{0})}$, $c_{0}$ as in Theorem 1, $j=\lceil\log(d)\rceil$, $h\geq 1$, we have that
$\mathbb{P}(L^{(j)}\geq K\mid X_{0}=x_{0},W_{0}=w_{0})\geq 1-\epsilon\,.$
Notice the validity of the bound in Corollary 1 for any starting value $X_{0}$ and innovation $W_{0}$. This allows to prove the next main theorem, where we bound with high probability the number of parallel iterations required by the Online Picard algorithm to simulate a given number of steps of RWM with $K=\mathcal{O}(\sqrt{d})$ parallel processors.
Theorem 2.
Suppose Assumptions 2-3 holds with $h\geq 1$. Then, for all $\delta>0$, $n\in\mathbb{N}$, $K\leq\sqrt{d/(8c_{0})}$, $c_{0}$ as in Theorem 1, we have that $\mathbb{P}(L^{(j)}>n)\geq 1-\delta$ with
$j=\lceil\log(d)\max\{4n/K,2\log(1/\delta)\}\rceil=\mathcal{O}\left(\frac{n}{K}\right)\,.$
By Theorem 2, the Online Picard algorithm achieves the optimal speedup relative to the sequential algorithm for $K$ up to $\mathcal{O}(\sqrt{d})$, since it simulates $n$ steps of the sequential algorithm with $\mathcal{O}(n/K)$ parallel iterations and $K$ parallel processors. Notice that the convergence of the Picard map does not deteriorate with the condition number $L/m$ of the target.
Combining Theorem 2 with available explicit bounds on the mixing times of RWM (Andrieu et al., 2024) gives the following complexity result.
Corollary 2 (Parallel round complexity of Online Picard algorithm).
Suppose Assumptions 2-3 holds with $mI_{d}\preccurlyeq\nabla^{2}V$, $h=1$ and $X_{0}\sim\mathcal{N}(x^{\star},L^{-1}I_{d})$, where $x^{\star}$ is the minimizer of $V$. Then, for any $\epsilon>0$, the Online Picard algorithm with $K=\lfloor\sqrt{d/(8c_{0})}\rfloor$, $c_{0}$ as in Theorem 1, outputs a random variable $X$, with $\|\mathcal{L}(X)-\pi\|_{\mathrm{TV}}\leq\epsilon$ after
$J=\mathcal{O}\left(\frac{L}{m}\sqrt{d}\,\mathrm{polylog}(\epsilon^{-1})\right)$
parallel iterations, where $\mathcal{L}(X)$ denotes the law of $X$ and $\|P-Q\|_{\mathrm{TV}}=\sup_{A}|P(A)-Q(A)|$ for two probability distributions $P$ and $Q$.
Corollary 2 implies that the Online Picard algorithm with $K=\mathcal{O}(\sqrt{d})$ achieves a Parallel round complexity of $\mathcal{O}(\sqrt{d})$.
3.4 Contraction of the Picard map: comparison between first- and zeroth-order case
In this section we compare the contraction of the Picard maps applied to RWM to the one obtained in the Unadjusted Langevin algorithm (ULA) case (Anari et al., 2024). The comparison illustrates how the first-order unadjusted case, where the map $x\mapsto f(x,W)$ is smooth, differs from the zeroth-order case of RWM, where $x\mapsto f(x,W)$ is piecewise constant.
Denote by $\Phi^{\mathrm{ULA}}\colon\mathcal{X}^{K+1}\times\mathcal{W}^{K}\mapsto\mathcal{X}^{K+1}$ the Picard map applied to the ULA Markov chain, defined as in (1) with
$f^{\mathrm{ULA}}(X,W)=-\frac{\xi^{2}}{2}\nabla V(X)+\xi W,\quad W\sim\mathcal{N}(0,I_{d})$
for a tuning parameter $\xi>0$. Under Assumption 2, for all $x,y\in\mathcal{X}^{K+1},$ with $x_{0}=y_{0}$, all $w\in\mathcal{W}^{K}$, and for all $i\leq K$, we have
$\|\Phi_{i}^{\mathrm{ULA}}(x,w)-\Phi_{i}^{\mathrm{ULA}}(y,w))\|$ $=\frac{\xi^{2}}{2}\|\sum_{\ell=0}^{i-1}(\nabla V(x_{\ell})-\nabla V(y_{\ell}))\|\leq\frac{\xi^{2}L}{2}\sum_{\ell=0}^{i-1}\|x_{\ell}-y_{\ell}\|,$
while Lemmas 5-6 in the supplement material imply that, under Assumptions 2-3,
$\mathbb{E}[\|\Phi^{\mathrm{RWM}}_{i}(x,W)-\Phi^{\mathrm{RWM}}_{i}(y,W)\|]$ $\leq\sqrt{\mathbb{E}[\|\Phi^{\mathrm{RWM}}_{i}(x,W)-\Phi^{\mathrm{RWM}}_{i}(y,W)\|^{2}]}$
$\leq\sqrt{15\xi^{2}d\sum_{\ell=1}^{i-1}\mathbb{P}(f(x_{\ell},W_{\ell})\neq f(y_{\ell},W_{\ell}))+\delta(d)}$
$\leq\sqrt{15\xi^{3}Ldc_{1}\sum_{\ell=1}^{i-1}\|x_{\ell}-y_{\ell}\|+\frac{1}{\xi Lc_{1}}\delta(d)}$ (10)
where $W\sim\nu^{\otimes K}$, with marginals as in (8), $c_{1}=(\sqrt{\frac{2}{\pi}}+\frac{\gamma\xi\sqrt{Ld}}{2})$ and $\delta(d)$ as in Theorem 1. Note that the contraction of the Picard map occurs pointwise (for each innovation $w$) for ULA and in expectation for RWM. More crucially, the RWM contraction has a slower rate, mainly due to the the square root function in (10).
3.5 Behavior in the tails
The convergence bounds shown in Section 3.3 apply to any initialization $X_{0}$. Here we show that the convergence is faster when the chain is in the tails of the distribution.
Proposition 2.
Suppose Assumptions 2-3 hold, $\ell I_{d}\preccurlyeq\nabla^{2}V$ and $X^{(0)}_{\ell}=x_{0}$ for all $\ell\geq 0$. Then, for all $j\geq 0$ and $K>0$, we have
$\lim_{|x_{0}|\to\infty}\mathbb{P}\left(L^{(j)}=jK\right)=1.$
Proposition 2 implies that, as the initial point goes further in tails, the convergence of the Picard map becomes faster (up to converging in one iteration). This is partly related to classical scaling limits obtained in the transient phase for RWM (Christensen et al., 2005), which show that that RWM in the tails has a deterministic behavior. In our context, Proposition 2 suggests that the Picard map converges faster in the transient phase, which we do observe in the numerical simulations of Section 6.3.
4 Metropolis within Gibbs
We now extend our convergence results to Metropolis within Gibbs (MwG) Markov chains. Our main motivation to consider MwG is that it displays better empirical performances in Section 6.
4.1 Metropolis within Gibbs kernels
For a given orthonormal basis $\{o_{0},o_{1},\dots,o_{d-1}\}$ spanning $\mathbb{R}^{d}$ (e.g. the standard basis), the (deterministic-scan) Metropolis within Gibbs (MwG) chain on $\mathbb{R}^{d}$ (Geman and Geman, 1984) performs iteratively a one-dimensional Metropolis-Hasting update in the subspace spanned by $o_{i}$, for $i=0,\dots,d-1$. The corresponding chain $(X_{i})_{i=0,1,\dots}$ can be treated as a time-inhomogeneous Markov chain, where every block of $d$ coordinates of the Markov chain is iteratively defined as
[…]
with innovations $W_{i}=(U_{i},\,Z_{i})$, where $Z_{0},\dots,Z_{d-1}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\nu$ are one dimensional symmetric random variables, $U_{0},\dots,U_{d-1}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Unif}([0,1])$, independently of $Z_{0},\dots,Z_{d-1}$ and
[…]
with $B$ as in (8).
4.2 Analysis
If we denote by $\mathrm{Haar}(d)$ the Haar measure on the orthogonal group in dimension $d$, our theoretical results hold with the following assumption.
Assumption 4.
$X_{0},X_{1},\dots$ is a MwG chain on $\mathcal{X}=\mathbb{R}^{d}$ with $(o_{0},o_{1},\dots,o_{d-1})\sim\mathrm{Haar}(d)$ and $Z=h/\sqrt{Ld}(2P-1)S$, $P\sim\mathrm{Ber}(1/2)$ and $S\sim\chi(d)$, where $\mathrm{Ber}(p)$ and $\chi(d)$ are respectively the Bernoulli distribution with parameter $p$ and the Chi distribution with $d$ degrees of freedom.
Theorem 3.
Under Assumptions 2 and 4,
  • the statement of Theorem 1 holds with $c_{0}$ and $\delta(d)$ replaced by $c_{0}=2h^{4}(\sqrt{\frac{2}{\pi}}+\frac{h\gamma}{2})^{2}$, $\delta(d)=11\exp(-d/10)$;
  • the statement of Theorem 2 holds with $h>5$, $c_{0}$ as above;
  • the statement of Proposition 4 below holds with $c_{0}$ as above and $d\geq-10\log(\epsilon r/33)$.
Theorem 3 shows that the analysis of the Picard map derived in Section 3.3 can be extended to MwG. However, for MwG, it is not possible to derive the parallel round complexity (as in Corollary 2) since, to the best of our knowledge, tight complexity bounds for MwG under a feasible start remain unknown, with the only available result known to us being a bound on the spectral gap provided in Ascolani et al. (2024, Corollary 7.4).
The next proposition shows that the Online Picard algorithm, when applied to MwG for isotropic Gaussian targets, achieves optimal speedup equal to $K$ with $K$ parallel processors, since each application of the Picard map converges instantaneously to its fixed point. This might be related to the greater speedups of MwG relative to RWM for well conditioned targets empirically observed in Section 6.
Proposition 3 (MwG with isotropic Gaussian targets).
Suppose Assumption 4 holds and $V(x)=\|x-\mu\|^{2}/(2\sigma^{2})$, $\mu\in\mathbb{R}^{d},\sigma^{2}>0$. For every $K\leq d$ and $X_{0}\in\mathcal{X}$, we have $L^{(j)}=jK$.
Note that the result in Proposition 3 is deterministic, i.e. it holds for every sequence $w\in\mathcal{W}^{\infty}$.
5 Approximate Picard
By Theorem 1 the probability of correctly guessing each individual increment for the Picard map remains small up to $i=\mathcal{O}(d)$ steps. However, when accumulating all these small probabilities across $i$, the first incorrect guess happens much earlier, at $i=\mathcal{O}(\sqrt{d})$, see Corollary 1. This creates a bottleneck and does not allow us to leverage effectively more than $K=\mathcal{O}(\sqrt{d})$ parallel processors, see Theorem 2.
This motivated us to propose an Approximate Online Picard algorithm, which can leverage effectively up to $K=\mathcal{O}(d)$ processors, but whose output does not match anymore the output of the recursion in (1) and for which the limiting distribution of the resulting Markov chain has a small bias relative to $\pi$.
5.1 Approximate Online Picard algorithm
By (4), each iteration of the Online Picard algorithm advances up to the position of the first error of the Picard recursion. Here we relax this condition by tolerating a fixed (and small) percentage $r\in[0,1]$ of mistakes. We thus define the Approximate Online Picard algorithm exactly as in (3) and (4), but replacing $(L^{(j)},U^{(j)})_{j=0}^{\infty}$ with $(L_{r}^{(j)},U_{r}^{(j)})_{j=0}^{\infty}$ defined as $L^{(0)}_{r}=0$, $U_{r}^{(j)}=L_{r}^{(j)}+K$ and
$L_{r}^{(j+1)}$ $=\sup\{L_{r}^{(j)}<i\leq U_{r}^{(j)}\colon\mathcal{A}_{\ell}^{(j)}\leq r\text{ for all }L_{r}^{(j)}<\ell\leq i\}$ $j\geq 0\,,$
where
$\mathcal{A}_{\ell}^{(j)}=\frac{|\{s\in\{L_{r}^{(j)},L_{r}^{(j)}+1,\dots,\ell-1\}\colon f(X^{(j+1)}_{s},W_{s})\neq f(X^{(j)}_{s},W_{s})\}|}{\ell-L_{r}^{(j)}}$
is the proportion of mistakes made by the $j$-th Picard recursion in $\{L_{r}^{(j)}+1,\dots,\ell\}$, see also the pseudocode in Appendix C. Here $r$ acts as a tolerance parameter which controls the approximation error, and the case $r=0$ coincides with the (exact) Online Picard algorithm. By construction $L_{r}^{(j)}$ is non-decreasing w.r.t. $r$, and thus increasing $r$ reduces the runtime. Interestingly, for any $r>0$, the Approximate Online Picard algorithm is able to exploit up to $K=\mathcal{O}(d)$ parallel processors effectively, as shown in the following Proposition.
Proposition 4.
Under Assumptions 2-3, for all $\epsilon>0$, $r\in(0,1]$, $d\geq\log(5/(\epsilon r))$, $j\geq 1.5\log(3/(\epsilon r))$ and $K\leq\epsilon rd/(3c_{0})$, we have
$\mathbb{P}\left(\frac{1}{K}\sum_{i=0}^{K-1}\mathbf{1}\left(f(X_{i}^{(j)},W_{i})\neq f(X_{i},W_{i})\right)\leq r\right)>1-\epsilon.$
Proposition 4 shows that the fraction of incorrect guesses can be controlled with high probability up to $K=\mathcal{O}(d)$ processors and with only $j=\mathcal{O}(1)$ parallel iterations. This suggests that the Approximate Online Picard algorithm with $r>0$ and $K=\mathcal{O}(d)$ reduces the parallel run-time of the Online Picard algorithm by a factor $\mathcal{O}(\sqrt{d})$, thus converging in $j=\mathcal{O}(1)$ parallel iterations, as numerically verified in Section 6.
However, for $r>0$ the output no longer matches exactly the output of the sequential algorithm. Intuitively, we expect the error introduced to be small for small values of $r$, since the increments are coherent with the ones of the original RwM at least $(1-r)$-fraction of the times, so that, the Markov chain at those steps follows the correct transition kernel in (1). Also, the size of the other increments is bounded by the size of $Z$, which has Gaussian tails. The error of the Approximate Picard algorithm on the invariant distribution is numerically assessed in Section 6 for tolerance levels $r$ ranging from 0% to 20%, while a rigorous analysis of the error introduced by taking $r>0$ is outside the scope of this work.
6 Numerical simulations
We assess numerically the performance of the (Approximate) Online Picard algorithm applied to RWM and MwG Markov chains for high-dimensional regression models and the SIR epidemic model. The former simulations allow to test performances systematically and assess the tightness of the bounds developed in Section 3. The latter experiment illustrates the breadth of our results in a more sophisticated Bayesian problem where the posterior density is not log-concave and is only piecewise continuous. In this setting, state-of-the-art gradient-based methods such as Langevin and Hamiltonian Monte Carlo (Roberts and Tweedie, 1996a; Neal et al., 2011) cannot be directly used and zeroth-order (gradient-free) methods become a natural choice.
6.1 Algorithmic specification and performance metrics
In all numerical simulations, we set $\xi=c^{\mathrm{(RWM)}}/\sqrt{d}$ for RWM, and innovation $\tilde{Z}=\mathcal{N}(0,c^{\mathrm{(MwG)}})$ for MwG (with standard basis). The stepsizes $c^{\mathrm{(RWM)}}$ and $c^{\mathrm{(MwG)}}$ are tuned to obtain an average Metropolis acceptance-rejection of approximately 23.4% and 40%, respectively, as prescribed by the literature (Gelman et al., 1997, 1996). The code to reproduce the simulations is available at https://github.com/SebaGraz/ParallelMH.git.
We define the empirical speedup relative to sequential implementations as the average number of sequential steps simulated per parallel round, i.e.
$\hat{G}=L^{(J)}/J\,,$
where $J$ is the total number of iterations of the (Approximate) Online Picard algorithm. We can interpret $\hat{G}$ as the multiplicative speedup obtained by parallel computing when ignoring overheads due to parallel execution.
We evaluate the bias introduced with the Approximate Online Picard algorithms by computing the standardized moments:
$\mathcal{M}_{r}=\sqrt{\frac{1}{d}\sum_{i=1}^{d}\frac{(\hat{\mu}^{(r)}_{i}-\mu_{i}(\pi))^{2}}{\sigma_{i}^{2}(\pi)}},\qquad\mathcal{E}_{r}=\sqrt{\frac{1}{d}\sum_{i=1}^{d}\left(\frac{\hat{\sigma}_{i}^{(r)}-\sigma_{i}(\pi)}{\sigma_{i}(\pi)}\right)^{2}}$ (12)
where $\hat{\mu}^{(r)}_{i}$ is the sample mean and $\hat{\sigma}_{i}^{(r)}$ is the sample standard deviation of the $i$th element estimated through the Markov chain $X_{0:N}$ simulated with the Approximate Picard algorithm with tolerance level $r$ while $\mu_{i}(\pi)$ and $\sigma_{i}(\pi)$ are the mean and standard deviation of the $i$th-marginal posterior distribution (obtained with a very long MCMC run, when not analytically tractable).
6.2 High-dimensional regressions
We consider three high-dimensional regression models:
  • E1
[…]
  • E2
[…]
  • E3
[…]
Figure 3 and Figure 4 summarize the results. The left panels display $\hat{G}$ as a function of $d$ (on a log-log scale), when $K=d$. The growth matches the one predicted by the theory of Section 3, i.e. $\mathcal{O}(\sqrt{d})$ for Online Picard and $\mathcal{O}(d)$ for the approximate version. The middle panels display $\hat{G}$ as a function of $K$ for fixed $d$ ($d=500$ for linear regression, $d=200$ otherwise). Again, the results align with the theory, with a linear speedup up to $K=\sqrt{d}$ (blue vertical dashed line) followed by a slower increase until $K=d$ and no increase afterwards. The right panel displays $\mathcal{M}_{r}$ and $\mathcal{E}_{r}$, defined in (12), as a function of $r$, for $K=d$ and $d=200$. Interestingly, for RWM, increasing $r$ tends to inflate the variance of the resulting approximate samples, in a way analogous to the bias observed in unadjusted Langevin algorithms (Roberts and Tweedie, 1996b; Dalalyan and Riou-Durand, 2020), while this is not the case for MwG, which appears to be significantly more robust to larger values of $r$. For these experiments, we set $N=10^{5}$ for the linear regression and $N=4\times 10^{5}$ for the logistic and Poisson regressions.
6.3 Convergence in the tails
We numerically assess the convergence of the Picard map for RWM chains started in the tails of the target distribution. Proposition 2 shows instantaneous convergence for log-concave targets in the limit for $|X_{0}|\to\infty$. Figure 5 numerically shows the convergence of $L^{(1)}$ to $K=200$ for initializations $X_{0}\sim\text{Unif}(x\in\mathbb{R}^{d}\colon\|x-x^{\star}\|=s)$, as $s$ gets larger, where $x^{\star}$ is the posterior mode and the target is given by the logistic model E2 in Section 6.2 with $d=200$ and a standard Gaussian prior. In all experiments, the tuning parameter $h$ was kept fixed and tuned in a preliminary run with the algorithm in stationarity. The results are averaged over 10 simulations.
Figure 3: Performance of Online Picard algorithm ($\bar{X}$) and its approximate versions ($\bar{X}_{r},\,r=0\%,\dots,20\%$) applied to RWM and MwG ($\bar{X}_{MwG}$), with target being the linear regression model E1. Left panel: average speedup $\hat{G}$ ($y$-axis, on a log-scale) with $K=d$, $N=10^{4}$ and $d=10^{2},\dots,10^{3}$ ($x$-axis, on a log-scale). Dashed lines $d\mapsto\sqrt{d}$ (blue) and $d\mapsto 3d/20$ (red) for reference. Middle panel: average speedup $\hat{G}$ ($y$-axis, on a log-scale) for $d=500$, $N=10^{4}$ and $K=2,\dots,10^{3}$ ($x$-axis, on a log scale). Vertical dash lines for $K=\sqrt{d}$ and $K=d$ for reference. Right panel: $\mathcal{M}$ and $\mathcal{E}$ as in (12) for the Approximate Picard algorithms with tolerance ($x$-axis) ranging from $0\%$ (exact Picard algorithm) to $20\%$ for $d=10^{2},\,K=d,\,N=10^{5}$.
Figure 4: Same as Figure 3 for the logistic regression model E2 (top panels) and the Poisson regression model E3 (bottom panels), with $N=10^{4},K=d$ (left panels), $N=10^{4},d=200$ (middle panels), $d=200,K=d,N=4\times 10^{5}$ (right panels).
Figure 5: $L^{(1)}$ ($y$-axis) for $X_{0}\sim\text{Unif}(x\in\mathbb{R}^{d}\colon\|x-x^{\star}\|=s)$, $s$ ranging from 0 to 2000 ($x$-axis) and target given by logistic regression E2 with $d=K=200$.
6.4 SIR epidemic model
Let $\{Y(t)\in\{S,I,R\}^{M},t\geq 0\}$ be an infection process in a population of size $M$. Each coordinate $Y_{i}(t)$ takes values
$y_{i}(t)$ =$=\begin{cases}S&\text{if $i$ is \emph{susceptible} at time $t$}\\ I&\text{if $i$ is \emph{infected} at time $t$}\\ R&\text{if $i$ is \emph{removed} at time $t$}\\ \end{cases},$ $i=1,2,\dots,M.$
Each coordinate process $t\to Y_{i}(t)$ is allowed to change state in the following direction: $S\to I\to R$. In particular the coordinate process changes its state from $S$ to $I$ according to an inhomogeneous Poisson process with rate $t\to\beta_{0}\mathcal{I}_{t}$ where $\beta_{0}>0$ and $\mathcal{I}_{t}=\sum_{i=1}^{M}\mathbf{1}(y_{i}(t)=I)$ is the total number of infected individuals at time $t$. The transition from $I$ to $R$ of each individual occurs according to an exponential random variable with rate equal to $\gamma_{0}>0$. The parameters $\beta_{0}$ and $\gamma_{0}$ are fundamental as they jointly determine the reproduction number $R_{0}=M\beta_{0}/\gamma_{0}$. See Baily (1975) for a thorough mathematical description of this model and its extensions.
Let $d\leq M$ be the total number of individuals that have been infected during the epidemic. We assume that the epidemic has originated with 1 initial infected individual and has been observed until the epidemic is extincted, that is, after the last infected infected individual has been removed.
[…]
$\exp\left(-\beta A(x,x^{\circ})+\sum_{\begin{subarray}{c}i=1\\ i\neq k(x)\end{subarray}}^{d}\log(\beta\mathcal{I}_{x_{i}-})\right)\\ \left(\prod_{i=1}^{d}\gamma\exp(-\gamma(x^{\circ}_{i}-x_{i}))\mathbf{1}(x_{i}\leq x^{\circ}_{i})\right).$
Here $k=\arg\min(x)$ is the label of the first infected individual, $t-=\lim_{s\uparrow t}s$ the left-hand limit and
$A(x,x^{\circ})=\int_{x_{k}}^{T}\mathcal{I}_{t}\mathcal{S}_{t}\mathrm{d}t=\sum_{j=1}^{d}\left((M-d)(x_{j}^{\circ}-x_{j})+\sum_{i=1}^{d}(x_{j}^{\circ}\wedge x_{i}-x_{j}\wedge x_{i})\right),$ (13)
with $T=\max(x^{\circ}_{1},x^{\circ}_{2},\dots,x^{\circ}_{d})$ and $\mathcal{S}_{t}=\sum_{i=1}^{M}\mathbf{1}(y_{i}(t)=S)$, is the total number of person-to-person units of infectious pressure exerted during the course of the epidemic, see (Neal and Roberts, 2005, Section 5) for details of the likelihood and posterior measure in the set-up considered here.
We set two independent Gamma priors for $\gamma$ and $\beta$ with prior parameters $(\nu_{\beta},\lambda_{\beta})$ and $(\nu_{\gamma},\lambda_{\gamma})$ respectively. We then integrate $\gamma,\beta$ out the posterior and get the marginal posterior density of the latent space of infection time:
$\pi(x)\propto\left(\prod_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{d}\mathcal{I}_{x_{i}-}\right)\left(\lambda_{\beta}+A(x,x^{\circ})\right)^{-(d+\nu_{\beta}-1)}\\ \left(\lambda_{\gamma}+\sum_{i=1}^{d}(x_{i}^{\circ}-x_{i})\right)^{-(d+\nu_{\gamma})}\left(\prod_{i=1}^{d}\mathbf{1}(x_{i}\leq x^{\circ}_{i})\right).$ (14)
This posterior density is only piecewise continuous, with regions of discontinuity in correspondence to changes of the order of the combined vector of infection and removal times $(x,x^{\circ})$. Once $x$ is simulated, we can obtain posterior samples of $\gamma$ and $\beta$ generating them from the conditional distributions
[…]
where $B(x,x^{\circ})=\sum_{i=1}^{d}(x^{\circ}_{i}-x_{i})$ and $A(x,x^{\circ})$ as in (13).
We simulate synthetic data and adopt the same experimental setting of Neal and Roberts (2005, Section 5.2), that is we set $M=200$, $\beta=0.001$ and $\gamma=0.15$. With this setting, the reproduction number is equal to $R=4/3$. The forward model generates $d=98$ infected individuals before the epidemic was extincted.
We set weak uninformative priors for $\beta$ and $\gamma$ with parameters $\lambda_{\beta}=\lambda_{\beta}=0.001$ and $\nu_{\beta}=\nu_{\beta}=1$ and run the Online-Picard applied to both RMW and MwG with target density as in (14) and for $N=10^{6}$ steps and with $K=\lfloor\sqrt{d}\rfloor=9$ parallel processors. The chain is initialized out of stationarity, at $X_{0,i}\sim x^{\circ}_{i}-\text{Exp}(0.05)$ (verifying that the posterior evaluated at this initial value is non-zero, that is $\pi(X_{0})\neq 0$). Moreover, we perform a second and computationally more intensive numerical experiment with $M=400,\,R=8/3,\,d=372,\,N=5\times 10^{6},\,K=\lfloor\sqrt{d}\rfloor=19$.
Table 1 shows the results obtained, where $\mathcal{M}_{0},\mathcal{E}_{0}$ were computed as in (12) relative to the $d$ dimensional latent observation $x$ with Markov chains with burn-in equal to $N/2$, while the true means and variances where estimated with a long run ($N=10^{7}$ for the first numerical experiment and $N=5\times 10^{7}$ for the second one, burn-in equal to $N/5$ in both cases) of RWM, started at the true values $x$. Figure 6 shows the functionals $B(X,x^{\circ})$ and $A(X,x^{\circ})$ as in (15), evaluated at every 10 steps of the RWM chain, with time index displayed with a yellow-blue gradient-color. The right panel, shows the corresponding values of $(\beta,\gamma)$, simulated from the conditional posteriors in (15).
Table 1: Online Picard applied to SIR model with $d=98(K=9,\,N=10^{6}$) and $d=372\,(K=19,\,N=5\times 10^{6})$.
Markov Chain $\hat{G}$ $\mathcal{M}_{0}$ $\mathcal{E}_{0}$
RWM $d=98$ 4.31 0.24 0.20
MwG $d=98$ 4.64 0.15 0.17
RWM $d=372$ 5.34 0.27 0.21
MwG $d=372$ 9.58 0.12 0.09
Figure 6: $(A(X_{i},x^{\circ}),B(X_{i},x^{\circ}))$ (left panel) and independent samples of $(\gamma,\beta)\mid(X_{10\,i},x^{\circ})$ (right panel), from $i=0$ (yellow) to $i=10^{5}$ (blue) and $X_{i}$ being the $i$th step of the RWM. The regions $R\geq 1$ is displayed in sky-blue. The true values from which we simulated the data are shown with red points.
7 Extensions
We outline here some extensions of this work, which are left for future research:
  • The Online Picard recursion (3) sets $X^{(j+1)}_{i}$ for $i\geq U^{(j)}$ to a constant value equal to the last coordinate of the Picard map $\Phi$. While this choice is simple and reasonable, it does not exploit the intrinsic dynamics of the underlying Markov chain and the noise variables beyond the time index $U^{(j)}$. More elaborate recursions are possible, combining Picard maps with inexpensive predictive functions that estimate the values of the Markov chain at those steps where the Picard map is not been applied yet. An example of such schemes is provided by the (stochastic) Parareal algorithm (Lions et al., 2001; Pentland et al., 2022).
  • By analyzing the family of approximate measures $\pi_{r}$ for $r>0$ in Section 5.1, it may be possible to establish the parallel round complexity of the Approximate Online Picard algorithm. Our numerical results suggest that this algorithm can achieve a parallel round complexity of order $\mathcal{O}(1)$ using $\mathcal{O}(d)$ parallel processors.
  • The proof of Proposition 1 shows that $L^{(j+1)}=L^{(j)}+G^{(j)}$ with $G^{(j)}=G(\bar{X}^{(j)},\bar{W}^{(j)})$ and $G$ a deterministic function. Assuming that $(\bar{X}^{(j)},\bar{W}^{(j)})_{j\in\mathbb{N}}$ has a unique invariant measure $\rho$, then we can interpret $\mathbb{E}_{\rho}(G(\bar{X},\bar{W}))$ as the average speedup relative to the sequential algorithm in stationarity and $N/\mathbb{E}_{\rho}(G(\bar{X},\bar{W}))$ is the expected number of parallel iterations of the Online Picard algorithm for simulating the Markov chain $X_{0:N}$. This suggest that trying to characterize explicitly the invariant distribution $\rho$ might be a promising way to improve the analysis of the Online Picard algorithm, e.g. going beyond classical log-concavity or obtaining more explicit expressions.
Acknowledgments
Both authors acknowledge support from the European Union (ERC), through the Starting grant ‘PrSc-HDBayLe’, project number 101076564.
Appendix A Proofs for Random Walk Metropolis
A.1 Background lemmas
We first state some background lemmas that will be helpful to streamline the proofs of the main results.
Lemma 1.
Let $X\sim\mathcal{E}(\lambda)$, where $\mathcal{E}(\lambda)$ denotes the exponential distribution with rate $\lambda$. Then for any $a,b\in\mathbb{R}$ we have $\mathbb{P}(X\in[a\wedge b,a\vee b])\leq\mathbb{P}(X\leq a\vee b-a\wedge b)$, where $a\wedge b$ and $a\vee b$ denote, respectively, the minimum and maximum between $a$ and $b$.
Proof.
Without loss of generality, let $a\leq b$. For the cases $a,b\leq 0$ and $a\leq 0,b\geq 0$, Lemma 1 is trivial, while for $a,b\geq 0$, it follows from the fact that the density function of an exponential random variable is non-increasing. ∎
Lemma 2.
$Z\sim\mathcal{N}(0,\sigma^{2}I_{d})$ with $\sigma\geq 0$. Then, for every $v\in\mathbb{R}^{d}$, we have $\mathbb{E}[|\langle v,Z\rangle|]=\|v\|\sqrt{\frac{2}{\pi}}\sigma.$
Proof.
By rotational invariance of $\mathcal{N}(0,\sigma^{2}I_{d})$, we have $\mathbb{E}[|\langle v,Z\rangle|=\|v\|\mathbb{E}[|Z_{1}|]$, where $|Z_{1}|$ follows a one-dimensional folded normal distribution and has mean $\sqrt{\frac{2}{\pi}}\sigma$. ∎
Lemma 3.
Let $Z\sim\mathcal{N}(0,\sigma^{2}I_{d})$ and $M$ be a $d\times d$ random matrix independent of $Z$. Then $\mathbb{E}[Z^{\top}MZ]=\sigma^{2}\mathrm{Tr}(\mathbb{E}[M])$.
Proof.
$\mathbb{E}[Z^{\top}MZ]=\sum_{i,j}\mathbb{E}[Z_{i}Z_{j}]\mathbb{E}[M_{i,j}]=\sum_{i}\mathbb{E}[Z_{i}^{2}]\mathbb{E}[M_{i,i}]=\sigma^{2}\mathrm{Tr}(\mathbb{E}[M])$. ∎
Lemma 4.
For any non-negative random variable $Z$ and constant $c>0$, we have
$\mathbb{E}[Z\mathbf{1}(Z>c)]=\int_{c}^{\infty}\mathbb{P}(Z\geq u)]\mathrm{d}u+c\mathbb{P}(Z>c).$
Proof.
Follows from
$\mathbb{E}[Z\mathbf{1}(Z>c)]-c\mathbb{P}(Z>c)$ $=\mathbb{E}[\int_{0}^{Z}\mathbf{1}(u>c)\mathrm{d}u]=\mathbb{E}[\int_{c}^{\infty}\mathbf{1}(u<Z)\mathrm{d}u]$
$=\int_{c}^{\infty}\mathbb{E}[\mathbf{1}(u<Z)\mathrm{d}u]=\int_{c}^{\infty}\mathbb{P}(Z\geq u)\mathrm{d}u\,,$
where we used Fubini-Tonelli’s Theorem in the third equality. ∎
A.2 Contraction of the Picard map: key lemmas
The following three lemmas are at the core of the analysis of the contraction of the Picard map $\Phi$, as well as of the proofs of the main results.
Lemma 5.
Under Assumptions 2-3, for every $x,y\in\mathcal{X}$,
$\mathbb{P}(f(x,W)\neq f(y,W))\leq\frac{hL^{1/2}}{d^{1/2}}\left(\sqrt{\frac{2}{\pi}}+\frac{h\gamma}{2}\right)\|x-y\|,\qquad W\sim\nu.$
Proof.
By (8), under Assumptions 2-3, we have that
$\mathbb{P}(f(x,W)\neq f(y,W))=\mathbb{P}(B(x,U,Z)\neq B(y,U,Z))$
with $U\sim\text{Unif}([0,1])$ and $Z\sim\mathcal{N}(0,h^{2}/(Ld)I_{d})$. Since $E=-\log(U)\sim\mathcal{E}(1)$, by Lemma 1 we have
$\mathbb{P}(B(x,U,Z)$ $\neq B(y,U,Z)\mid Z)$
$=\left|\mathbb{P}\left(E<V(x+Z)-V(x)\right)-\mathbb{P}\left(E<V(y+Z)-V(y)\right)\right|$
$\leq 1-\exp(-|\Delta|)$
where $\Delta=V(x+Z)-V(x)-(V(y+Z)-V(y))$. Thus
$\mathbb{P}(B(x,U,Z)\neq B(y,U,Z))$ $=\mathbb{E}[\mathbb{P}(B(x,U,Z)\neq B(y,U,Z)\mid Z)]$
$\leq\mathbb{E}[1-e^{-|\Delta|}]$
$\leq\mathbb{E}[|\Delta|]\,,$
since $1-e^{-x}\leq x,x\in\mathbb{R}$.
Since $V$ is twice-continuously differentiable we have
$\Delta=\underbrace{(\nabla V(y)-\nabla V(x))^{\top}Z}_{\Delta_{1}}+\underbrace{\frac{1}{2}Z^{\top}(\nabla^{2}V(x+tZ)-\nabla^{2}V(y+tZ))Z}_{\Delta_{2}}$
for some $t=t(x,y)\in[0,1]$. By the triangle inequality
$\mathbb{E}[|\Delta|]\leq\mathbb{E}[|\Delta_{1}|]+\mathbb{E}[|\Delta_{2}|].$
For the first term, by $Z\sim\mathcal{N}(0,c^{2}/(Ld)I_{d})$, Lemma 2 and the $L$-smoothness of $V$, we have
$\mathbb{E}[|\Delta_{1}|]=\sqrt{\frac{2}{\pi}}\frac{hL^{1/2}}{d^{1/2}}\|\nabla V(y)-\nabla V(x)\|\leq\sqrt{\frac{2}{\pi}}\frac{h}{(Ld)^{1/2}}\|x-y\|^{2}.$
For the second term, by the strong Hessian-Lipschitz condition in Assumption 2 and Lemma 3, we have
$\mathbb{E}[|\Delta_{2}|]$ $=\frac{1}{2}\mathbb{E}[|Z^{\top}(\nabla^{2}V(x+tZ)-\nabla^{2}V(y+tZ))Z|]$
$\leq\frac{1}{2}\mathbb{E}[|Z^{\top}MZ]\|x-y\|$
$=\frac{h^{2}}{2}\frac{\mathrm{Tr}(M)}{Ld}\|x-y\|$
$\leq\frac{h^{2}}{2}\frac{\gamma L^{3/2}d^{1/2}}{Ld}\|x-y\|$
$=\frac{h^{2}\gamma L^{1/2}}{2d^{1/2}}\|x-y\|\,.$
Hence, we have
$\mathbb{E}[|\Delta|]$ $\leq\mathbb{E}[|\Delta_{1}|]+\mathbb{E}[|\Delta_{2}|]\leq\frac{hL^{1/2}}{d^{1/2}}\left(\sqrt{\frac{2}{\pi}}+\frac{h\gamma}{2}\right)\|x-y\|$
as desired. ∎
Lemma 6.
Under Assumption 3 and for all $x,y\in\mathcal{X}^{K+1}$ with $x_{0}=y_{0}$, $w_{0}\in\mathcal{W}$ and $1<i\leq d$,
$\mathbb{E}[\|\Phi_{i}(x,W)-\Phi_{i}(y,W)\|^{2}\mid W_{0}=w_{0}]$ $\leq\frac{15h^{2}}{L}\sum_{\ell=1}^{i-1}\left(\mathbb{P}(f(x_{\ell},W_{\ell})\neq f(y_{\ell},W_{\ell}))+\delta(d)\right),$ (16)
with $W_{1:K-1}\sim\nu^{\otimes(K-1)}$, $\nu^{\otimes p}$ being product measure of $\nu$ repeated $p$ times, and $\delta(t)$ as in Theorem 1. The inequality (16) holds also when $x$ and $y$ are replaced by random variables $X,Y$ on $\mathcal{X}^{K+1}$, which can depend on $W$.
The proof of Lemma 6 relies on the following bounds.
Lemma 7.
Let $W\sim\hbox{Wishart}(I_{i},d)$ be a Wishart matrix on $\mathbb{R}^{i\times i}$ with $d$ degrees of freedom and identity scale matrix, and $Y$ be a random vector on $\mathbb{R}^{i}$ defined on the same probability space, with $\|Y\|^{2}\leq i$ almost surely. Then
$\mathbb{E}[Y^{T}WY]\leq 15\mathbb{E}[\|Y\|^{2}]+i25\exp(-3d/2)\,.$
Proof.
By Wainwright [2019, Theorem 6.1], for all $t\geq 0,\,i\leq d$,
$\mathbb{P}\left(\lambda_{\max}(W)>(1+\sqrt{i/d}+t)^{2}\right)\leq e^{-t^{2}d/2}$
for its maximum eigenvalue. Combining this result and Lemma 4, we have
$\mathbb{E}[\lambda_{\max}(W)\mathbf{1}(\lambda_{\max}(W)\geq 15)]$ $=\int_{15}^{\infty}\mathbb{P}(\lambda_{\max}(W)\geq u)\,\mathrm{d}u+15\mathbb{P}(\lambda_{\max}(W)>15)$
$\leq\int_{15}^{\infty}e^{-(\sqrt{u}-2)^{2}d/2}\mathrm{d}u+15e^{-d(\sqrt{15}-2)^{2}/2}$
$\leq\int_{15}^{\infty}e^{-ud/10}\mathrm{d}u+15e^{-3d/2}$
$=\frac{10}{d}\exp(-3d/2)+15\exp(-3d/2)$
$\leq 25\exp(-3d/2)$
where we used $(\sqrt{u}-2)^{2}>u/5$ for $u\geq 15$. Finally,
$\mathbb{E}[Y^{T}WY]$ $\leq\mathbb{E}[\lambda_{max}(W)\mathbf{1}(\lambda_{max}(W)\leq 15)\|Y\|^{2}]+i\mathbb{E}[\lambda_{max}(W)\mathbf{1}(\lambda_{max}(W)>15)])$
$\leq 15\left(\mathbb{E}[\|Y\|^{2}]+i\delta(d)\right)$
with $\delta(d)=\frac{5}{3}\exp(-3d/2)$. ∎
Proof of Lemma 6.
By the definition of $\Phi_{i}$ in (2) and $f$ in (8), we have
$\Phi_{i}(x,W)-\Phi_{i}(y,W)$ $=\sum_{\ell=1}^{i-1}\Delta B_{\ell}Z_{\ell}$ (17)
[…]
$\|\sum_{\ell=1}^{i-1}\Delta B_{\ell}Z_{\ell}\|^{2}=\sum_{\ell,j=1}^{i-1}\Delta B_{\ell}\Delta B_{j}\langle Z_{\ell},Z_{j}\rangle\stackrel{{\scriptstyle d}}{{=}}\frac{h^{2}}{L}\Delta B^{T}W\Delta B$ (18)
with $\Delta B=(\Delta B_{1},\dots,\Delta B_{i-1})^{T}$ and $W\sim\hbox{Wishart}(I_{i-1},d)$. Here =$\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution. Combining (17)-(18) and using Lemma 7 we have
$\mathbb{E}[\|\Phi_{i}(x,W)-\Phi_{i}(y,W)\|^{2}\mid W_{0}=w_{0}]=\frac{h^{2}}{L}\mathbb{E}[\Delta B^{T}W\Delta B]\leq\frac{15h^{2}}{L}(\mathbb{E}[\|\Delta B\|^{2}]+i\delta(d))$
with $\delta(d)=\frac{5}{3}\exp(-3d/2)$. The result follow by $\mathbb{E}[\|\Delta B\|^{2}]=\sum_{\ell=1}^{i-1}\mathbb{P}(B(x_{\ell},U_{\ell},Z_{\ell})\neq B(y_{\ell},U_{\ell},Z_{\ell}))$. ∎
Fix $i\in\{0,\dots,d\}$ and define
$A^{(j)}$ $=\max_{\ell\leq i}\mathbb{P}(f(X^{(j)}_{\ell},W_{\ell})\neq f(X_{\ell},W_{\ell})),$ $j\in\{0,1,\dots\}\,.$
Lemma 8.
Under the assumptions of Theorem 1, we have
$A^{(j+1)}\leq\sqrt{c_{0}\frac{i}{d}\left(A^{(j)}+\delta(d)\right)}$
with $c_{0}$, $\delta(d)$ as in Theorem 1.
Proof.
By Lemma 5 and Jensen’s inequality we have
$\mathbb{P}(f(X^{(j+1)}_{i},W_{i})\neq f(X_{i},W_{i}))$ $=\mathbb{E}[\mathbb{P}(f(X^{(j+1)}_{i},W_{i})\neq f(X_{i},W_{i})\mid X^{(j+1)}_{i},X_{i})]$
$\leq\frac{hL^{1/2}}{d^{1/2}}\left(\sqrt{\frac{2}{\pi}}+\frac{h\gamma}{2}\right)\mathbb{E}[\|X^{(j+1)}_{i}-X_{i}\|]$
$\leq\frac{hL^{1/2}}{d^{1/2}}\left(\sqrt{\frac{2}{\pi}}+\frac{h\gamma}{2}\right)\sqrt{\mathbb{E}[\|X^{(j+1)}_{i}-X_{i}\|^{2}]}\,.$
By Lemma 6 we have that
$\mathbb{E}[\|X^{(j+1)}_{i}-X_{i}\|^{2}]$ $=\mathbb{E}[\|\Phi_{i}(X^{(j)},W)-\Phi_{i}(X,W)\|^{2}]$
$\leq\frac{15h^{2}}{L}\sum_{\ell=1}^{i-1}\left(\mathbb{P}(f(X^{(j)}_{\ell},W_{\ell})\neq f(X_{\ell},W_{\ell}))+\delta(d)\right)$
$\leq\frac{15h^{2}}{L}i\left(A^{(j)}+\delta(d)\right)\,.$
Combining the above gives
$\mathbb{P}(f(X^{(j+1)}_{i},W_{i})\neq f(X_{i},W_{i}))$ $\leq\sqrt{15}h^{2}\left(\sqrt{\frac{2}{\pi}}+\frac{h\gamma}{2}\right)\sqrt{\frac{i}{d}\left(A^{(j)}+\delta(d)\right)}\,,$
which implies the desired result after noting that $\max_{\ell\leq i}$ is a non decreasing operator in $i$. ∎
A.3 Proof of Theorem 1
Lemma 9.
Let $(a_{j})_{j=0,1,\dots}$ be a non-negative sequence satisfying $a_{0}=1$ and $a_{j+1}\leq b\sqrt{a_{j}+\epsilon}$ for some fixed $b,\epsilon>0$ and all $j\geq 0$. Then $a_{j}\leq b^{2}+\epsilon+2^{-j}$ for all $j\geq 0$.
Proof.
Let $g(a)=b\sqrt{a+\epsilon}$. By monotonicity of $g$, we can assume without loss of generality that $a_{0}=1$ and $a_{j+1}=g(a_{j})$ for all $j\geq 1$. The function $g$ is concave and its unique fixed point on $[0,\infty)$ satisfying $a^{*}=g(a^{*})$ is
$a^{*}=\frac{b^{2}+\sqrt{b^{4}+4\epsilon b^{2}}}{2}\in(b^{2},b^{2}+\epsilon)\,.$
Indeed, imposing $g(a)^{2}=a^{2}$ gives $a^{2}-b^{2}a-b^{2}\epsilon=0$, whose solutions are $(b^{2}+\sqrt{b^{4}+4\epsilon b^{2}})/2$ and $(b^{2}-\sqrt{b^{4}+4\epsilon b^{2}})/2$, and the latter is negative. Also, $(b^{2}+\sqrt{b^{4}+4\epsilon b^{2}})/2\geq(b^{2}+\sqrt{b^{4}})/2=b^{2}$ and, by concavity of $x\mapsto\sqrt{x}$, $\sqrt{b^{4}+4\epsilon b^{2}}\leq b^{2}+\frac{1}{2\sqrt{b^{4}}}4\epsilon b^{2}=b^{2}+2\epsilon$ which implies $a^{*}=(b^{2}+\sqrt{b^{4}+4\epsilon b^{2}})/2\leq(b^{2}+b^{2}+2\epsilon)/2=b^{2}+\epsilon$.
If $a_{0}=1\leq a^{*}$ then $a_{j}\leq a^{*}\leq b^{2}+\epsilon\leq b^{2}+\epsilon+2^{-j}$ for all $j\geq 0$, as desired. Consider now $a_{0}=1\geq a^{*}$. Since $g^{\prime}$ is strictly decreasing on $[0,\infty)$, for all $a\geq a^{*}$ we have $20<g^{\prime}(a)\leq g^{\prime}(a^{*})\leq g^{\prime}(b^{2}-\epsilon)=1/$ 2. Thus, $g$ is a strict contraction on $[a^{*},\infty)$ with Lipschitz constant smaller than $1/2$ and, by Banach’s fixed point theorem, we have $|a_{j}-a^{*}|\leq 2^{-j}|a_{0}-a^{*}|$ for all $j\geq 0$. By $a_{0}=1\geq a^{*}$, we have $|a_{0}-a^{*}|\leq a_{0}=1$ and thus we deduce $a_{j}\leq a^{*}+2^{-j}\leq b^{2}+\epsilon+2^{-j}$ as desired. ∎
Proof of Theorem 1.
. Follows directly by Lemma 8 and Lemma 9 with $b=\sqrt{c_{0}i/d}$, $\epsilon=\delta(d)$, and $c_{0},\delta(d)$ as in Theorem 1. ∎
A.4 Proof of Corollary 1
By (4) we have
$\mathbb{P}(L^{(j)}<K)$ $=1-\mathbb{P}\left(\bigcap_{i=0}^{K-1}f(X^{(j)}_{i},W_{i})=f(X_{i},W_{i})\right)$
$=\mathbb{P}\left(\bigcup_{i=0}^{K-1}f(X^{(j)}_{i},W_{i})\neq f(X_{i},W_{i})\right)\,.$
Applying the union bound and the bound in Theorem 1, for $K\leq b\sqrt{d},b>0$, it follows
$\mathbb{P}(L^{(j)}<K)$ $\leq\sum_{i=0}^{K-1}\mathbb{P}(f(X^{(j)}_{i},W_{i})\neq f(X_{i},W_{i}))\leq b\sqrt{d}\left(\frac{c_{0}b}{\sqrt{d}}+\delta(d)+2^{-j}\right)\,.$
This is less than $\epsilon$ if $c_{0}b^{2}=\frac{\epsilon}{2}$ and $b\sqrt{d}(\delta(d)+2^{-j})\leq\epsilon/2$. The first condition is achieved by setting $b=\frac{\epsilon}{\sqrt{2c_{0}}}$. With this choice, the latter condition is always satisfied for $h\geq 1,\,d\geq 1$ since $(h,d)\mapsto\sqrt{\frac{d}{2c_{0}}}(\delta(d)+2^{-j})-\frac{1}{2}$ is negative on $[1,\infty)^{2}$.
A.5 Proof of Theorem 2
For two real-valued random variables $X\sim\mu,\,Y\sim\nu$, we write $X\succeq Y$ if $\mu([x,\infty))\geq\nu([x,\infty))$ for all $x\in(-\infty,\infty)$.
Proposition 5.
Under Assumption 2-3, $h\geq 1$, for all $\epsilon>0$, $K\leq\sqrt{\epsilon^{2}d/(2c_{0})}$, $x_{0}\in\mathcal{X}$, any $j\in\mathbb{N}$ and $p=\lceil\log(d)\rceil$, we have
$(L^{(jp)}\mid X_{0}=x_{0})\succeq\sum_{i=1}^{j}P_{i},$
with $P_{i}\overset{i.i.d.}{\sim}(1-\epsilon)\delta_{K}(\cdot)+\epsilon\delta_{0}(\cdot)$.
The proof of Proposition 5 requires the following two lemmas. For any $i\in\mathbb{N}$, define $\tilde{G}_{i}:=\sum_{\ell=0}^{p-1}G^{(pi+\ell)}$, with $(G^{(i)})_{i=0,1,\dots}$ defined in (7).
Lemma 10.
Suppose that, for some random variable $A$, $\tilde{G}_{i}\mid\tilde{G}_{0:(i-1)},A\succeq P_{i}$. Then $\tilde{G}_{i}\mid\tilde{G}_{0:(i-1)}\succeq P_{i}$.
Proof.
By the law of total probability we have
$\mathbb{P}(\tilde{G}_{i}\geq t\mid\tilde{G}_{0:(i-1)})=\mathbb{E}_{A}[\mathbb{P}(\tilde{G}_{i}\geq t\mid G_{0:(i-1)},A)]\geq\mathbb{E}_{A}[\mathbb{P}(P_{i}\geq t)]=\mathbb{P}(P_{i}\geq t).$
Lemma 11.
Suppose $\tilde{G}_{i}\mid\tilde{G}_{0:(i-1)}\succeq P_{i}$ for all $i\geq 0$. Then $\sum_{i=0}^{j}\tilde{G}_{i}\succeq\sum_{i=0}^{j}P_{i}$ for all $j\geq 0$.
Proof.
Let $\hat{G}_{j}:=\sum_{i=0}^{j}\tilde{G}_{i}$, $\hat{P}_{j}:=\sum_{i=0}^{j}P_{i}$. Then $\hat{G}_{0}=\tilde{G}_{0}\succeq P_{0}=\hat{P}_{0}$. By induction, if $\hat{G}_{j}\succeq\hat{P}_{j}$, for some $j$, then we have
$F_{\hat{G}_{j+1}}(t)=F_{\tilde{G}_{j+1}+\hat{G}_{j}}(t)$ $=\int F_{\tilde{G}_{j+1}\mid\hat{G}_{j}}(t-u)\mathrm{d}F_{\hat{G}_{j}}(u)$
$\leq\int F_{P_{i+1}}(t-u)\mathrm{d}F_{\hat{G}_{j}}(u)$
$\leq\int F_{P_{i+1}}(t-u)\mathrm{d}F_{\hat{P}_{j}}(u)=F_{\hat{P}_{j+1}}(t)$
where $F_{X}(t)$ denotes the distribution function of $X$ evaluated at $t$ and the last inequality follows since $u\mapsto F_{P_{i+1}}(t-u)$ is decreasing. ∎
Proof of Proposition 5.
As shown in the proof of Proposition 1, we have $L^{(j+1)}=\sum_{i=0}^{j}G^{(i)}$. Let $\tilde{L}_{i}:=L^{(pi)}=\sum_{\ell=0}^{i-1}\tilde{G}_{\ell}$. Note that $\{\tilde{L}_{i}=l\}\in\sigma(X_{0},W_{0:l})$ where $\sigma(X)$ denotes the sigma-algebra generated by the random variable $X$.
Let $A=(X_{0},W_{0:\tilde{L}_{i}},\tilde{L}_{i})$. Then $\tilde{G}_{0:(i-1)}\in\sigma(A)$,
$\mathbb{P}(\tilde{G}_{i}\geq\ell\mid A,\tilde{G}_{0:(i-1)})=\mathbb{P}(\tilde{G}_{i}\geq\ell\mid A)$
and, for any $a=(x_{0},w_{0:l},l)$, $\epsilon>0$ and $K\leq\epsilon\sqrt{d/(2c_{0})}$ we have
$\mathbb{P}(\tilde{G}_{i}\geq\ell\mid A=a)$ $=\mathbb{P}(\tilde{G}_{i}\geq\ell\mid X_{\tilde{L}_{i}}=x_{l},W_{\tilde{L}_{i}}=w_{l},\tilde{L}_{i}=l)$
$=\mathbb{P}(\tilde{G}_{0}\geq\ell\mid X_{\tilde{L}_{0}}=x_{l},W_{\tilde{L}_{0}}=w_{l})$
$=\mathbb{P}(L^{(p)}\geq\ell\mid X_{0}=x_{l},\,W_{0}=w_{l})$
$\geq\mathbb{P}(P_{i}\geq\ell)\,,$
where $x_{i}=x_{0}+\sum_{\ell=1}^{i-1}f(x_{\ell},w_{\ell}),\,i\geq 1$ and the last inequality follows from Corollary 1 and from $\tilde{G}_{i}\geq p$ almost surely, which holds by construction. Combining the above with Lemmas 10-11 yields the desire result. ∎
Proof of Theorem 2.
Fix $\epsilon=1/2$. By Proposition 5, for all $j\in\mathbb{N}$, $p=\lceil\log(d)\rceil$, $K\leq\sqrt{d/(8c_{0})}$, we have
$\mathbb{P}(L^{(pj)}<n\mid X_{0}=x)$ $=\mathbb{P}(\sum_{i=0}^{j-1}G^{(i,p)}<n\mid X_{0}=x)\leq\mathbb{P}\left(\sum_{i=1}^{j}P_{i}<n\right)\,$
with $(P_{i})_{i=1}^{j}$ as in Proposition 5, $\mu=\mathbb{E}[P_{1}]=(1-\epsilon)K=K/2$. By Hoeffding’s inequality, we have
$\mathbb{P}\left(\sum_{i=1}^{j}P_{i}<n\right)$ $=\mathbb{P}\left(\sum_{i=1}^{j}P_{i}-j\mu<-(j\mu-n)\right)$
$\leq\exp\left(-\frac{2(j\mu-n)^{2}}{jK^{2}}\right)$
$\leq\exp\left(-\frac{2(jK/2-n)^{2}}{jK^{2}}\right)=\exp\left(-j\left(1-\frac{2n}{jK}\right)\right)$
for any $j>2n/K$. In particular, if $j\geq\max\{\frac{4n}{K},2\log(1/\delta)\}$, then $j\left(1-\frac{2n}{jK}\right)\geq 2\log(1/\delta)\left(1-\frac{1}{2}\right)=\log(1/\delta)$ and the above exponential bound is less than $\delta$. ∎
A.6 Proof of Corollary 2
Let $(X_{n})_{n=0,1,2,\dots}$ be the RWM trajectory, i.e. the exact solution of the recursion in (1). Under Assumptions 2-3, $mI_{d}\preccurlyeq\nabla^{2}V$, $h=1$ and $X_{0}\sim\mathcal{N}(x^{\star},L^{-1}I_{d})$, we can apply Theorem 49 and Remark 50 of Andrieu et al. [2024] to deduce $\sqrt{\chi^{2}(\mathcal{L}(X_{n_{mix}}),\pi)}\leq\epsilon$ with $n_{mix}=\mathcal{O}(\frac{L}{m}d)$, where $n_{mix}$ depends on $\epsilon$, $d$, $L/m$ and is explicitly provided in Andrieu et al. [2024, Rmk.50]. Consider running the Online Picard algorithm with $K=\sqrt{d/(8c_{0})}$ for
$J=\lceil\log(d)\max\{4n_{mix}/K,2\log(2/\epsilon)\}$
parallel iterations and returning $Y=X^{(J)}_{\min\{n_{mix},L^{(J)}\}}$ as output. When $n_{mix}\leq L^{(J)}$, we have $Y=X^{(J)}_{n_{mix}}=X_{n_{mix}}$. By Theorem 2, $\mathbb{P}(Y=X_{n_{mix}})\geq\mathbb{P}(L^{(J)}\geq n_{mix})\geq 1-\epsilon/2$. Therefore
$\|\mathcal{L}(Y)-\pi\|_{TV}$ $\leq\mathbb{P}(Y\neq X_{n_{mix}})+TV(\mathcal{L}(X_{n_{mix}}),\pi)$
$\leq\epsilon/2+\frac{1}{2}\sqrt{\chi^{2}(\mathcal{L}(X_{n_{mix}}),\pi)}$
$\leq\epsilon/2+\epsilon/2=\epsilon\,.$
A.7 Proof of Proposition 2
For a sequence $x_{n}\to\infty$, let $A_{n}=V(x_{n}+Z)-V(x_{n})$, $B_{n}=V(Y_{n}+Z)-V(Y_{n})$, $Z\sim\mathcal{N}(0,1)$, where $Y_{n}=x_{n}+N_{n}$ and for a sequence of random variables $(N_{n})_{n=1,2,\dots}$ on $\mathbb{R}^{d}$ with $\mathbb{E}[\|N_{n}\|^{2}]\leq C$.
Lemma 12.
Under Assumption 2 with $mI_{d}\preccurlyeq\nabla^{2}V$, we have
$\mathbb{E}[|A_{n}-B_{n}|]$ $\leq 2L(C+d),$ for all $\hbox{for all }n=1,2,\dots\,,$ (19)
$\limsup_{n}\mathbb{P}(|A_{n}|\leq c)$ $=0$ for $\hbox{for all }n=1,2,\dots\,,\,c\in(0,\infty).$ (20)
Proof.
Let $A=V(x+z)-V(x)$, $B=V(y+z)-V(y)$, for some $x,y,z\in\mathcal{X}$. By $L$-smoothness of $V$, we have that, for some $t_{1},t_{2}\in(0,1)$ depending on $x,y,z$,
$|A-B|$ $=|\langle\nabla V(x+t_{1}z)-\nabla V(y+t_{2}z),z\rangle|$
$\leq\|\nabla V(x+t_{1}z)-\nabla V(y+t_{2}z)\|\|z\|$
$\leq L(\|x-y\|+\|z\|)\|z\|$
$\leq 2L\|x-y\|^{2}+\|z\|^{2}.$
Thus,
$\mathbb{E}[|A_{n}-B_{n}|]$ $\leq 2L\mathbb{E}[\|x_{n}-Y_{n}\|^{2}+\|Z\|^{2}]$
$=2L(\mathbb{E}[\|N_{n}\|^{2}]+\mathbb{E}[\|Z\|^{2}])$
$\leq 2L(C+d)$
proving (19). Also, again by $L$-smoothness of $V$,
$|A_{n}|\geq|\langle\nabla V(x_{n}),Z\rangle|-L\|Z\|^{2}/2\,$
thus, for all $c\in(0,\infty)$
$\mathbb{P}(|A_{n}|\leq c)\leq\mathbb{P}(\|Z\|^{2}\geq\frac{2}{L}(|\langle\nabla V(x_{n}),Z\rangle|-c)).$ (21)
By rotational invariance $Z$, the right hand-side of (21) depends on $\nabla V(x_{n})$ only through its norm. By $m$-convexity of $V$, $\|\nabla V(x_{n})\|\to\infty$ as $|x_{n}|\to\infty$ and, consequently, the right hand-side of (21) goes to 0, proving the desired result in (20). ∎
Lemma 13.
Under Assumption 2 with $mI_{d}\preccurlyeq\nabla^{2}V$, we have
$\limsup_{n\to\infty}\mathbb{E}[|\exp(-\max\{A_{n},0\})-\exp(-\max\{B_{n},0\})|]=0\,.$
Proof.
For a fixed $c\in(0,\infty)$, we have that
$\mathbb{E}[|\exp(-\max\{A_{n},0\})-\exp(-\max\{B_{n},0\})|]\\ \leq e^{-c}+\mathbb{P}(|A_{n}|\leq c\cap|B_{n}|\leq c)+\mathbb{P}(|A_{n}-B_{n}|>2c).$
By Lemma 12, $\limsup_{n\to\infty}\mathbb{P}(A_{n}>c\cap B_{n}>c)=0$ and
$\mathbb{P}(|A_{n}-B_{n}|>2c)\leq\frac{\mathbb{E}[|A_{n}-B_{n}|]}{2c}\leq\frac{L(C+d)}{c}.$
Combining these results and since $c$ was arbitrary, we have
$\limsup_{n\to\infty}\mathbb{E}[|\exp(-\max\{A_{n},0\})-\exp(-\max\{B,0\})|]$ $\leq\inf_{c}\{e^{-c}+\frac{L(C+d)}{c}\}=0\,.$
Proof Proposition 2.
Here we prove $\lim_{n\to\infty}\mathbb{P}(L^{(j)}=jK)=1$ for $j=1$. The case $j\geq 1$ follows by recursion. By (8), we have
$1-\mathbb{P}(L^{(1)}=K)$ $=1-\mathbb{P}\left(\bigcap_{i=0}^{K}X^{(1)}_{i}=X_{i}\right)$
$=\mathbb{P}\left(\bigcup_{i=0}^{K}\left(X^{(1)}_{i}\neq X_{i}\right)\right)$
$=\mathbb{P}\left(\bigcup_{i=0}^{K-1}\left(B(X^{(0)}_{i},W_{i})\neq B(X_{i},W_{i})\right)\right)$
$\leq\sum_{i=0}^{K-1}\mathbb{P}\left(B(x_{n},W_{i})\neq B(X_{i},W_{i})\right).$
Hence, $\lim_{n\to\infty}\mathbb{P}(L^{(1)}=K)=1$, if $\lim_{n\to\infty}\mathbb{P}\left(B(x_{n},W_{i})\neq B(X_{i},W_{i})\right)=0$ for all $i=0,1,\dots,K-1$.
For a fixed $i$, note that $X_{i}=x_{n}+N_{n}$ where $N_{n}=\sum_{\ell=0}^{i-1}B(X_{\ell},U_{\ell},Z_{\ell})Z_{\ell}$ and
$\mathbb{E}[\|N_{n}\|^{2}]$ $=\mathbb{E}[\|\sum_{\ell=0}^{i-1}B(X_{\ell},U_{\ell},Z_{\ell})Z_{\ell}\|^{2}]$
$\leq\sum_{\ell=0}^{i-1}\mathbb{E}[\|Z_{\ell}\|^{2}]+\sum_{\begin{subarray}{c}\ell,m=0\\ \ell\neq m\end{subarray}}^{i-1}\mathbb{E}[\|Z_{\ell}\|]\mathbb{E}[\|Z_{m}\|]\leq C,$
with $C=i^{2}d$. Then, for $E=-\log(U_{i})\sim\mathrm{Exp}(1)$,
$\mathbb{P}(B(x_{n},U_{i},Z_{i})\neq B(X_{i},U_{i},Z_{i}))$
$=\mathbb{E}[\mathbb{P}\left(B(x_{n},U_{i},Z_{i})\neq B(x_{n}+N_{n},U_{i},Z_{i})\mid U_{0:i-1},\,Z_{0:i}\right)]$
$=\mathbb{E}[\left|\mathbb{P}\left(E<V(x_{n}+Z)-V(x_{n})\right)-\mathbb{P}\left(E<V(x_{n}+N_{n}+Z)-V(x_{n}+N_{n})\right)\right|]$
$=\mathbb{E}[|e^{-\max(V(x_{n}+Z)-V(x_{n}),0)}-e^{-\max(V(x_{n}+N_{n}+Z)-V(x_{n}+N_{n}),0)}|]$
$=\mathbb{E}[|e^{-\max(A_{n},0)}-e^{-\max(B_{n},0)}|]$
which goes to 0 as $n\to\infty$ by Lemma 13. ∎
A.8 Proof of Proposition 4
By Markov inequality and the bound in Theorem 1, we have for $K\leq\lfloor bd\rfloor$, $b=\epsilon r/(3c_{0})$,
$\mathbb{P}\Big{(}\frac{1}{K}\sum_{i=0}^{K-1}\mathbf{1}\big{(}f(X_{i}^{(j)},W_{i})\neq f(X_{i},W_{i})\big{)}\leq r\Big{)}$ $\leq\frac{1}{Kr}\sum_{i=0}^{K-1}\mathbb{E}\left[\mathbf{1}\left(f(X_{i}^{(j)},W_{i})\neq f(X_{i},W_{i})\right)\right]$
$=\frac{\sum_{i=0}^{\lfloor bd\rfloor}\mathbb{P}(f(X^{(j)}_{i},W_{i})\neq f(X_{i},W_{i}))}{\lfloor bd\rfloor r}$
$\leq\frac{c_{0}b+\delta(d)+2^{-j}}{r}\leq\epsilon$
for $d\geq-2\log(\epsilon r/5)/3$, $j\geq-\log(\epsilon r/3)/\log(2)$.
Appendix B Proofs for Metropolis within Gibbs
B.1 Proof of Theorem 3
Theorem 3 essentially shows that most of the results derived in Section 3.3 holds (with different constants) also for Metropolis within Gibbs (MwG) Markov chains, i.e. when replacing Assumption 3 with Assumption 4.
Lemmas 5-6 are the building blocks of all the results of Section 3.3. Thus, we start by providing the analogue of these Lemmas for MwG.
Lemma 14.
Under Assumptions 2 and 4, the statement of Lemma 5 holds.
To prove Lemma 14, we first need the following proposition.
Proposition 6.
Let $W^{\mathrm{MwG}}$ and $W^{\mathrm{RWM}}$ be the innovations of MwG and RWM satisfying respectively Assumption 4 and Assumption 2. Then, for every $x\in\mathcal{X}$
$\mathcal{L}(f_{i}(x,W^{\mathrm{MwG}}))$ $=\mathcal{L}(f(x,W^{\mathrm{RWM}}))$ for $\hbox{for }i=0,1,\dots,d-1,$
with $f$ as in (8) and $f_{i}$ as in (11).
Proof.
By (8) and (11), Proposition 6 follows if $\mathcal{L}(o_{i}Z^{\mathrm{MwG}})=\mathcal{L}(Z^{\mathrm{RWM}})$ for all $i$, where $Z^{\mathrm{MwG}}=h/\sqrt{Ld}(2P-1)S$, $P\sim\mathrm{Ber}(1/2)$, $S\sim\chi(d)$, $Z^{\mathrm{RWM}}\sim\mathcal{N}(0,I_{d}h^{2}/(Ld))$, $(o_{0},o_{1},\dots,o_{d-1})\sim\mathrm{Haar}(d)$. By rotational invariance of $(o_{0},o_{1},\dots,o_{d-1})$ and $Z^{\mathrm{RWM}}$, it is enough to show that
$\mathcal{L}(\|o_{i}Z^{\mathrm{MwG}}\|)=\mathcal{L}(\|Z^{\mathrm{RWM}}\|)=\mathcal{L}(h/\sqrt{Ld}\chi(d)).$
Proof of Lemma 14.
By Proposition 6, Lemma 5 follows also under Assumptions 2-4, with exactly the same constants and proof provided in Section A.2. ∎
Next, we provide the analogue of Lemma 6 for MwG.
Lemma 15.
Under Assumption 4 and for all $x,y\in\mathcal{X}^{K+1}$, with $x_{0}=y_{0}$, $w_{0}\in\mathcal{W}$ and $1<i\leq d$
$\mathbb{E}[\|\Phi_{i}(x,W)-\Phi_{i}(y,W)\|^{2}\mid W_{0}=w]$ $\leq\frac{2h^{2}}{L}\sum_{\ell=1}^{i-1}\left(\mathbb{P}(f_{i}(x_{\ell},W_{\ell})\neq f_{i}(y_{\ell},W_{\ell}))+\delta(d)\right)$ (22)
with $\delta(d)=11\exp(-d/10)$ and where the expectation is relative to $(W_{1},W_{2},\dots,W_{K-1})$.
The proof of Lemma 15 relies on the following bounds.
Lemma 16.
Let $X$ and $Y$ be two random variables defined on the same probability space, with $X=\tilde{X}/d,\,\tilde{X}\sim\chi^{2}_{d}$ and $Y$ binary, i.e. $\mathbb{P}(Y\in\{0,1\})=1$. Then
$\mathbb{E}[XY]\leq 2\left(\mathbb{P}(Y=1)+11\exp(-d/10)\right)\,.$
Proof.
By Laurent and Massart [2000, Lemma 1], for every $t\geq 2$,
$\mathbb{P}(X>t)\leq e^{-td/20}.$
Then, Combining this result and Lemma 4, we have that
$\mathbb{E}[\mathbf{1}(X\geq 2)X]$ $=\int_{2}^{\infty}\mathbb{P}(X\geq u)\,\mathrm{d}u+2\mathbb{P}(X>2)$
$\leq\int_{2}^{\infty}e^{-ud/20}\mathrm{d}u+2e^{-d/10}$
$=\frac{20}{d}\exp(-d/10)+2\exp(-d/10)$
$\leq 22\exp(-d/10).$
Finally
$\mathbb{E}[YX]$ $\leq\mathbb{E}[YX\mathbf{1}(X\leq 2)]+\mathbb{E}[(X\mathbf{1}(X>2)]$
$\leq 2\left(\mathbb{E}[Y]+\delta(d)\right)$
with $\delta(d)=11\exp(-d/10)$. ∎
Proof of Lemma 15.
Under Assumptions 2 and 4,
$\Phi_{i}(x,W)-\Phi_{i}(y,W)$ $=\sum_{\ell=1}^{i-1}\Delta B_{\ell}Z_{\ell}o_{\ell}$
where $\Delta B_{\ell}=B(x_{\ell},U_{\ell},o_{\ell}Z_{\ell})-B(y_{\ell},U_{\ell},o_{\ell}Z_{\ell})$, $Z_{\ell}=h/\sqrt{Ld}(2P_{\ell}-1)S$, $U_{1},U_{2},\dots,U_{i-1}\overset{\mathrm{i.i.d}}{\sim}\text{Unif}([0,1])$, $P_{1},P_{2},\dots,P_{i-1}\overset{\mathrm{i.i.d}}{\sim}\mathrm{Ber}(1/2)$ and $S_{1},S_{2}\dots,S_{i-1}\overset{\mathrm{i.i.d}}{\sim}\chi(d)$.
Since $\langle o_{i},o_{j}\rangle=0$ almost surely for all $i\neq j$ and by Proposition 6, we have
$\|\sum_{\ell=1}^{i-1}\Delta B_{\ell}o_{\ell}Z_{\ell}\|^{2}=\sum_{\ell,j=1}^{i-1}\Delta B_{\ell}\Delta B_{j}\langle o_{\ell}Z_{\ell},o_{j}Z_{j}\rangle=\sum_{\ell=1}^{i-1}\Delta B_{\ell}^{2}\|o_{\ell}Z_{\ell}\|^{2}=\frac{h^{2}}{L}\sum_{\ell=1}^{i-1}\Delta B_{\ell}^{2}\tilde{Z}_{\ell}$
with $\tilde{Z}_{1},\,\tilde{Z}_{2},\dots,\tilde{Z}_{i-1}\overset{\mathrm{i.i.d}}{\sim}\chi^{2}(d)$. Thus, by Lemma 16, we have
$\mathbb{E}[\|\sum_{\ell=0}^{i-1}\Delta B_{\ell}o_{\ell}Z_{\ell}\|^{2}]$ $=\frac{h^{2}}{L}\sum_{\ell=1}^{i-1}\mathbb{E}[\Delta B_{\ell}^{2}\tilde{Z}]$
$\leq\frac{2h^{2}}{L}\sum_{\ell=1}^{i-1}\left(\mathbb{P}(B(x_{\ell},U_{\ell},o_{\ell}Z_{\ell})\neq B(y_{\ell},U_{\ell},o_{\ell}Z_{\ell}))+\delta(d)\right)$
with $\delta(d)=11\exp(-d/10)$. ∎
Proof of Theorem 3.
Under Assumptions 2 and 4, the statements of Lemma 8 and Theorem 1 hold with new constants
$c_{0}=2h^{4}\left(\sqrt{\frac{2}{\pi}}+\frac{h\gamma}{2}\right)^{2},\qquad\delta(d)=11\exp(-d/10).$ (23)
The proofs follow the same steps of the proof of Lemma 8 and Theorem 1 in Section A.3.
Similarly, Under Assumptions 2 and 4, the statements of Corollary 1 and Theorem 2 hold with constants $c_{0},\delta(d)$ as in (23), $K=\epsilon\sqrt{\frac{d}{2c_{0}}}$ and $h\geq 5$. The proofs follow the same steps of the proof of Corollary 1 in Section A.5 and the proof of Theorem 2 in Section A.4.
Finally, Under Assumptions 2 and 4 the statement of Proposition 4 holds with $c_{0}$ as in (23) and $d\geq-10\log(\epsilon r/33)$. ∎
B.2 Proof of Proposition 3
Let $\pi(x)\propto\exp(-\|x\|^{2}/(2\sigma^{2}))$ and let $X^{(0)}_{i}=X_{0}$ for $i=0,1,\dots,K$. We prove here that $L^{(1)}=K$, for all $X_{0}=x_{0}\in\mathcal{X}$, $K\leq d$. Proposition 3 then follows by recursion.
For MwG, $L^{(1)}{=}K$ if, for all $w=(w_{0},w_{1},\dots,w_{k-1})\in\mathcal{W}^{K}$,
$f_{i}(x_{0},w_{i})$ ${=}f_{i}(x_{i},w_{i})$ for all $\hbox{for all }0\leq i<K$ (24)
where $(x_{0},x_{1},x_{2}\dots,x_{K})$ is defined by the recursion $x_{i+1}=x_{i}+f_{i}(x_{i},w_{i})$, $f_{i}$ as in (11).
By (11), for all $u=(u_{0},u_{1},\dots,u_{K-1})\in[0,1]^{K}$, $(z_{0},z_{1},\dots,z_{K-1})\in\mathbb{R}^{K}$, the condition in (24) is equivalent to
$B(x_{0},u_{i},\hat{z}_{i})$ ${=}B(x_{i},u_{i},\hat{z}_{i})$ for all $\hbox{for all }0\leq i<K$ (25)
with $\hat{z}_{i}=o_{i}z_{i}$ and a given orthonormal basis $o_{0},o_{1},\dots,o_{d-1}$, $d\geq K$. Then, the left-hand side of (25) is
$B(x_{i},u_{i},\hat{z}_{i})$ $=B(x_{0}+\sum_{\ell=0}^{i-1}B(x_{\ell},u_{\ell},\hat{z}_{\ell})\hat{z}_{\ell},u_{i},\hat{z}_{i})$
$=\mathbf{1}\left(\frac{\pi(x_{0}+\sum_{\ell=0}^{i-1}B(x_{\ell},u_{\ell},\hat{z}_{\ell})\hat{z}_{\ell}+\hat{z}_{i})}{\pi(x_{0}+\sum_{\ell=0}^{i-1}B(x_{\ell},u_{\ell},\hat{z}_{\ell})\hat{z}_{\ell})}>u_{i}\right)$
$=\mathbf{1}\left(\frac{\|\hat{z}_{i}\|^{2}}{2\sigma^{2}}+\frac{\langle x_{0}+\sum_{\ell=0}^{i-1}B(x_{\ell},u_{\ell},\hat{z}_{\ell})\hat{z}_{\ell},\hat{z}_{i}\rangle}{\sigma^{2}}>u_{i}\right),$ (26)
while the right-hand side of (25) is
$B(x_{0},u_{i},z_{i})$ $=\mathbf{1}\left(\frac{\pi(x_{0}+\hat{z}_{i})}{\pi(x_{0})}>u_{i}\right)$
$=\mathbf{1}\left(\frac{\|\hat{z}_{i}\|^{2}}{2\sigma^{2}}+\frac{\langle x_{0},\hat{z}_{i}\rangle}{\sigma^{2}}>u_{i}\right).$ (27)
Because $o_{0},o_{1},\dots,o_{d-1}$ is an orthonormal basis on $\mathbb{R}^{d}$ we have that
$\langle\hat{z}_{\ell},\hat{z}_{i}\rangle$ $=z_{\ell}z_{i}\langle o_{\ell},o_{i}\rangle{=}0$ for all $\hbox{for all }i\neq\ell,$
hence, (26) and (27) are equal.
Appendix C Pseudocode
Algorithm 1 Online Picard algorithm
1:$X_{0}\in\mathcal{X}$, $N\in\mathbb{N}$, $K\in\mathbb{N}$.
2:Initialize $\bar{X}^{c}_{i}=\bar{X}_{i}\leftarrow X_{0}$, for $i=0,1,\dots,K$.
3:Set $L\leftarrow 0$
[…]
5:while $L\geq N$ do
6:    $\bar{X}\leftarrow\Phi(\bar{X}^{c},\bar{W})$
7:    $G\leftarrow\sup\{i\leq K\colon\bar{X}_{\ell}=\bar{X}^{c}_{\ell}\text{ for }0\leq\ell\leq i\}$
8:    $\bar{X}^{c}_{0}\leftarrow\bar{X}_{G}$
9:    for $i=0,1,\dots,K-1$ do
10:       if $i\leq K-G-1$ then
11:          $\bar{X}^{c}_{i+1}\leftarrow\bar{X}_{G+i+1}$
12:          $\bar{W}_{i}\leftarrow\bar{W}_{G+i}$
13:       else
14:          $\bar{X}^{c}_{i+1}\leftarrow\bar{X}_{K}$
15:          $\bar{W}_{i}\sim\nu$
16:       end if
17:    end for
18:    $L\leftarrow L+G$
19:end while
20:return $\bar{X}_{G-(L-N)}$
Our implementation of the Online Picard algorithm is based on Proposition 1 and is provided in Algorithm 1. Its output is the random variable $X_{N}$, for any $N\geq 0$ given as input. The Approximate Online Picard algorithm is identical to Algorithm 1 except for line 5, which is replaced by
$G\leftarrow\sup\{1\leq i\leq K\colon\bar{\mathcal{A}}_{\ell}\leq r\text{ for all }\ell\leq i\},$
where
$\bar{\mathcal{A}}_{\ell}=\frac{|\{0\leq s<\ell\colon f(\bar{X}_{s},\bar{W}_{s})\neq f(\bar{X}^{c}_{s},\bar{W}_{s})\}|}{\ell}\,.$